### v2

# Constrained free energy minimization for the design of thermal states and stabilizer thermodynamic systems

Abstract
A quantum thermodynamic system is described by a Hamiltonian and a list of conserved, non-commuting charges, and a fundamental goal is to determine the minimum energy of the system subject to constraints on the charges. Recently, [Liu et al., arXiv:2505.04514] proposed first- and second-order classical and hybrid quantum-classical algorithms for solving a dual chemical potential maximization problem, and they proved that these algorithms converge to global optima by means of gradient-ascent approaches. In this paper, we benchmark these algorithms on several problems of interest in thermodynamics, including one- and two-dimensional quantum Heisenberg models with nearest and next-to-nearest neighbor interactions and with the charges set to the total $x$, $y$, and $z$ magnetizations. We also offer an alternative compelling interpretation of these algorithms as methods for designing ground and thermal states of controllable Hamiltonians, with potential applications in molecular and material design. Furthermore, we introduce stabilizer thermodynamic systems as thermodynamic systems based on stabilizer codes, with the Hamiltonian constructed from a given code’s stabilizer operators and the charges constructed from the code’s logical operators. We benchmark the aforementioned algorithms on several examples of stabilizer thermodynamic systems, including those constructed from the one-to-three-qubit repetition code, the perfect one-to-five-qubit code, and the two-to-four-qubit error-detecting code. Finally, we observe that the aforementioned hybrid quantum-classical algorithms, when applied to stabilizer thermodynamic systems, can serve as alternative methods for encoding quantum information into stabilizer codes at a fixed temperature, and we provide an effective method for warm-starting these encoding algorithms whenever a single qubit is encoded into multiple physical qubits.
I Introduction
I.1 Background and motivation
Quantum computation has the potential to cause a paradigm shift in several areas of computer science. After the initial proposals of quantum speedups for factoring [1] and unstructured search [2], it was postulated that quantum computation could offer speedups for optimization, with applications to physics, combinatorial optimization, machine learning, and beyond (see, e.g., [3] for a recent review). While this possibility still remains the subject of ongoing research, concrete quantum algorithms for solving semi-definite optimization problems have been proposed starting nearly a decade ago [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Some of these algorithms have guaranteed runtimes [4, 5, 6, 7, 8, 10, 11, 16], while others are heuristic in nature [9, 12, 13, 14, 15], falling under the umbrella of variational quantum algorithms [17].
Liu et al. recently identified a concrete link between quantum thermodynamics and semi-definite optimization [16], having the benefit of bridging two previously unlinked foundational areas and bringing insights from quantum thermodynamics into the analysis of semi-definite optimization. As identified in [16], a special class of semi-definite optimization problems, known as constrained energy minimization problems, are relevant in quantum thermodynamics. In particular, these problems are defined in terms of a Hamiltonian and a list of conserved, non-commuting charges, with the latter imposing constraints on the state of the physical system. The goal then is to minimize the energy of the system, while respecting the imposed constraints. Generally speaking, determining the minimum energy of a quantum physical system is a primary goal in physics [18], typically being the first step employed in computing energetic properties of molecules and materials, and thus having applications in materials science [19], condensed-matter physics [20], and quantum chemistry [21]. Going beyond this application, and as discussed in this paper, constrained energy minimization has applications in the design of molecules and materials having desired properties.
Beyond establishing a link between quantum thermodynamics and semi-definite optimization, Ref. [16] proposed various classical and hybrid quantum-classical (HQC) algorithms for solving semi-definite programs, being motivated by physical intuition coming from quantum thermodynamics when applied to the special class of constrained energy minimization problems. These algorithms, hereafter referred to as the LMPW algorithms, have provable guarantees for their runtimes and are based on the observation that parameterized thermal states, also known as non-Abelian thermal states [22, 23, 24] or quantum Boltzmann machines [25, 26, 27], are optimal for solving constrained free-energy minimization problems, which closely approximate the original constrained energy minimization problems for sufficiently low temperatures. Indeed, Ref. [16] established that the optimization landscapes of dual chemical potential maximization problems are concave in the parameters of the thermal states, so that conventional optimization techniques like gradient ascent, stochastic gradient ascent, and their variants [28] are guaranteed to converge. The HQC algorithms are variational in nature, using a quantum computer for two purposes only, which include the preparation of parameterized thermal states and estimating expectations of observables (i.e., the Hamiltonian and non-commuting charges). A classical computer is then employed for storing and updating the parameters of the parameterized thermal states as the HQC algorithm progresses toward convergence. The LMPW HQC algorithms make use of first-derivative information to determine the next step in a search and can additionally incorporate second-derivative information to improve their convergence to globally optimal solutions.
Ref. [16] also proved that the HQC algorithms are sample efficient, meaning that they require a number of thermal-state preparations polynomial in the number of qubits in order to converge. However, for the HQC algorithms to have efficient overall runtimes, and not merely efficient sample complexity, the Hamiltonian and charges should be such that it is possible to efficiently prepare their corresponding low-temperature parameterized thermal states on a quantum computer. While there has been much theoretical progress on this topic in recent years [29, 30, 31, 32, 33, 34, 35, 36], it remains largely open to design quantum algorithms that can prepare the needed thermal states with temperature sufficiently low for accurately solving constrained energy minimization problems.
Let us also recall an observation from [16], that the LMPW HQC algorithms for solving constrained free energy minimization problems can alternatively be viewed as HQC algorithms for preparing non-Abelian thermal states of thermodynamic systems with conserved, non-commuting charges. This aspect of the LMPW HQC algorithms could be useful in fields like materials science, condensed matter physics, and quantum chemistry, because one expects such states to arise in these thermodynamic systems and, after preparing them, one could subsequently measure various observables of interest on the generated states, in order to determine their properties.
I.2 Summary of contributions
In this paper, we apply the LMPW algorithms to concrete examples of constrained energy minimization problems of interest in quantum thermodynamics, including the one- and two-dimensional Heisenberg models [37] with nearest and next-to-nearest neighbor interactions, and with the non-commuting charges set to the total magnetizations in the $x$, $y$, and $z$ directions. Such models describe interacting spin systems on a lattice and are essential for understanding magnetic materials [38].
As a first conceptual contribution, we offer an alternative compelling interpretation of the LMPW algorithms as methods for designing ground and thermal states of controllable Hamiltonians, such that these states satisfy desired properties (i.e., constraints on the expectation values of certain observables). This application of the algorithms could find use in the design of molecules and materials with desired properties.
As a key theoretical contribution, our paper also establishes a bridge between quantum error correction [39, 40, 41, 42, 43] and conserved, non-commuting charges in quantum thermodynamics, by observing that every stabilizer code [44, 45] can be understood as corresponding to a thermodynamic system with conserved, non-commuting charges. In particular, we set the Hamiltonian to be in correspondence with the stabilizer operators of the code, while the non-commuting charges are in correspondence with the logical operators of the code. We refer to such a thermodynamic system as a stabilizer thermodynamic system. Let us note that stabilizer Hamiltonians were proposed in [46] and have been investigated substantially since then (see, e.g., [47, 48, 49, 50, 51, 52, 53, 54, 55]). However, interpreting stabilizer and logical operators in terms of Hamiltonians and conserved, non-commuting charges, respectively, in a thermodynamic system has not been considered previously, to the best of our knowledge.
Moreover, we utilize this correspondence to demonstrate that the LMPW HQC algorithms can serve as thermodynamics-inspired HQC algorithms for encoding quantum information into quantum error-correcting codes, which is rather different from the standard approach that employs an encoding circuit to do so (see, e.g., [56, Chapter 4] and [57, Section 3.5]). The approach is instead based on constrained energy minimization and the fact that the LMPW HQC algorithms, as applied in this context, prepare low-temperature thermal states of stabilizer thermodynamic systems. We suspect that this approach will be useful in experimental scenarios for which it is accessible to prepare low-temperature thermal states of stabilizer thermodynamic systems.
As further examples of constrained energy minimization, and building on the aforementioned bridge, we also benchmark the LMPW algorithms on stabilizer thermodynamic systems constructed from the one-to-three-qubit repetition code [58], the perfect one-to-five-qubit code [59, 60], and the two-to-four qubit error-detecting code [61, 62]. For each case, we demonstrate that the LMPW HQC algorithms are capable of steering the state of the stabilizer thermodynamic system in such a way that any desired quantum state is encoded into the code’s logical qubits approximately, with the approximation improving as the temperature of the system decreases.
On the one hand, the numerical contributions of our paper can be viewed as complementing the theoretical developments of [16], by providing a comprehensive numerical study of the performance of the LMPW algorithms for constrained energy minimization problems. The value of this contribution is that it offers a practical assessment of the theoretical guarantees from [16] for various problems of interest in quantum thermodynamics. On the other hand, the first conceptual contribution of our paper, i.e., the design of ground and thermal states, could have far reaching applications for the use of quantum computers in molecular and material design. Additionally, the other theoretical contribution of our paper, i.e., establishing a bridge between quantum error correction and conserved, non-commuting charges in quantum thermodynamics via the notion of stabilizer thermodynamic systems, is conceptually enlightening and the related encoding algorithm could be useful as an alternative approach for stabilizing quantum information.
I.3 Overview of numerical simulations
Before proceeding with the details of our paper, let us first elaborate on some aspects of the numerical simulations that we conducted. Given that all of the constrained energy minimization problems we consider can be formulated as semi-definite programs (SDPs) [63, Section 4.6.2], the performance of the LMPW algorithms can be benchmarked against the values returned by standard SDP solvers. These solvers are based on the interior-point method [64, 65, 66] and are so reliable that we can view their outputs as essentially the true optimal values of the constrained energy minimization problems. We test out four variants of the LMPW algorithms, which include a first-order classical algorithm, a second-order classical algorithm, a first-order HQC algorithm, and a second-order HQC algorithm. Considering the classical algorithms in addition to the HQC algorithms is useful, as the classical algorithms can be understood as fully noiseless simulators of the HQC algorithms, such that the only kind of error involved when using them is the algorithmic error and not any other kind of error, e.g., shot noise arising from estimating expectations of observables. For this reason, we can benchmark the HQC algorithms against the classical algorithms, in addition to the true values generated by standard SDP solvers.
We conducted all of our simulations for small system sizes, between three to six qubits. These sizes are reasonable for providing a sense of the performance of the LMPW algorithms, while taking a reasonable amount of time for them to converge when conducted on readily available laptop computers. In order to speed up the second-order HQC algorithms, which require estimating the Hessian, we parallelized the implementation of them in order to use multiple processor cores. All of the software code needed to reproduce the results of our paper is publicly available as a zenodo repository https://zenodo.org/records/17285628. We leave it to future research to scale up the examples to much larger numbers of qubits, possibly by using tensor-network methods [67, 68, 69], and to conduct the simulations on actual quantum computers, incorporating techniques like error correction [41] and mitigation [70] in order to reduce the effects of noise.
I.3.1 Method for thermal state preparation
As mentioned in Section I.1, a critical subroutine used by the first- and second-order HQC algorithms is thermal state preparation, and there has been much development on this topic in recent years [29, 30, 31, 32, 33, 34, 35, 36]. However, in our paper, we do not make use of these more recent advanced methods. Instead, we employ a simple method for thermal state preparation, consisting of feeding the numerical values of the desired thermal state to Qiskit’s DensityMatrix routine [71] and employing its functionality for generic state preparation when given a description of the state. Although this method is elementary, it has value in demonstrating the performance of the algorithms under the assumption that a robust method for thermal state preparation exists. Indeed, the literature on thermal-state preparation algorithms is rather active and advances are being made rapidly, such that we expect continued progress in this direction during the near term. As such, any of these alternative thermal state preparation algorithms can be plugged in to our approach, and further study of the LMPW HQC algorithms’ performance can be conducted under them.
When using the aforementioned elementary method for thermal state preparation, it is essential to make use of the following identity, holding for all $b\in\mathbb{R}$ and every Hermitian matrix $A$, in order for this approach to be numerically stable, especially at low temperature $T$:
$\frac{\exp(A)}{\operatorname{Tr}[\exp(A)]}=\frac{\exp(A+bI)}{\operatorname{Tr}[\exp(A+bI)]}.$ (1)
For the specific case of the algorithms considered in our paper, we make the choices $A=-\frac{1}{T}(H-\mu\cdot Q)$ and $b=\frac{1}{T}\lambda_{\min}\!\left(H-\mu\cdot Q\right)$, so that the identity in (1) applies as follows to our case:
$\frac{\exp\!\left(-\frac{1}{T}(H-\mu\cdot Q)\right)}{\operatorname{Tr}\!\left[\exp\!\left(-\frac{1}{T}(H-\mu\cdot Q)\right)\right]}=\\ \frac{\exp\!\left(-\frac{1}{T}(H-\mu\cdot Q)+bI\right)}{\operatorname{Tr}\!\left[\exp\!\left(-\frac{1}{T}(H-\mu\cdot Q)+bI\right)\right]}$ (2)
In all of our numerical implementations, we make use of the right-hand side of (2) when encoding the numerical values of a thermal state at temperature $T$. Specifically, we evaluate the matrix $\frac{1}{T}(H-\mu\cdot Q)+bI$, apply the matrix exponential function, and then normalize the resulting matrix by its trace. Adding the specific value of $b$ given above has the effect of balancing out extremely low eigenvalues of $\frac{1}{T}(H-\mu\cdot Q)$, so that the result of the matrix exponential function is numerically stable (not doing so leads to numerical instabilities). See Sections II.1 and II.2.1 for more details of the thermal state in (2).
I.4 Paper organization
Our paper is organized as follows. Section II provides review material required for understanding the rest of the paper. In particular, Section II.1 reviews quantum thermodynamics in the presence of conserved, non-commuting charges, Section II.2 reviews the LMPW classical and HQC algorithms for constrained energy minimization, Section II.3 reviews quantum Heisenberg models, and Section II.4 reviews quantum error correction codes with a focus on stabilizer codes.
Our contributions start with Section III, where we offer an alternative interpretation of the LMPW algorithms as methods for designing Hamiltonians with ground and thermal states satisfying desired properties. In Section IV, we observe that every stabilizer code corresponds to a stabilizer thermodynamic system, in which the Hamiltonian is formed from the code’s stabilizer operators and the conserved, non-commuting charges are formed from the code’s logical operators. In Section V, we argue how the LMPW HQC algorithms for constrained energy minimization can be used to encode quantum information into stabilizer codes when applied to stabilizer thermodynamic systems. We then detail our simulation results in Section VI, exhibiting the performance of the LMPW classical and HQC algorithms on various quantum Heisenberg models and stabilizer thermodynamic systems. We finally conclude in Section VII with a summary of our findings and suggestions for future research.
II Review
II.1 Quantum thermodynamics with conserved, non-commuting charges
Recall that a qudit is a $d$-dimensional quantum system, where $d\in\mathbb{N}$, and a qubit is a qudit such that $d=2$. A thermodynamic system of $n\in\mathbb{N}$ qudits is described by a Hamiltonian $H$ and a tuple
$Q\coloneqq(Q_{1},\ldots,Q_{c})$ (3)
of non-commuting charges, where $c\in\mathbb{N}$. Each of these matrices is Hermitian and has dimension $d^{n}\times d^{n}$. The charges are conserved if each of them commutes with the Hamiltonian; i.e., if
$H,Q_{i}]=0\quad\text{for all}\quad i\in\left[c\right]\equiv\{1,\ldots,c\}.$ (4)
However, it is not required that the charges in $Q$ form a commuting set, and much of the recent interest in quantum thermodynamics regards the scenario in which the charges do not form a commuting set [22, 72, 73, 23, 74, 75, 24].
A charge $Q_{i}$ is extensive if it can be written in the following additive form:
$Q_{i}=\sum_{j=1}^{n}C_{i}^{(j)},$ (5)
where here and throughout we adopt the following shorthand:
$C_{i}^{(j)}\equiv I^{(1)}\otimes\cdots\otimes I^{(j-1)}\otimes C_{i}^{(j)}\otimes I^{(j+1)}\otimes\cdots\otimes I^{(n)},$ (6)
which leaves tensor products with identity operators implicit if not written explicitly. In (6), the superscript of each operator indicates the system on which the operator acts, $I^{(k)}$ denotes the $d\times d$ identity operator acting on system $k\in\left[n\right]$, and $C_{i}^{(j)}$ denotes a single-site charge operator acting on the $j$th system. In typical cases of extensive charges, we have that $C_{i}^{(j)}=C_{i}^{(k)}$ for all systems $j,k\in[n]$, so that the charge $Q_{i}$ is equal to the sum of the $i$th charges of each system.
Examples of conserved, non-commuting, and extensive charges that we consider in the quantum Heisenberg model (see Section II.3) include the total $X$, $Y$, and $Z$ magnetizations:
$X_{\operatorname{tot}}\coloneqq\sum_{j=1}^{n}X^{(j)},\quad Y_{\operatorname{tot}}\coloneqq\sum_{j=1}^{n}Y^{(j)},\quad Z_{\operatorname{tot}}\coloneqq\sum_{j=1}^{n}Z^{(j)},$ (7)
where $X^{(j)}$, $Y^{(j)}$, and $Z^{(j)}$ denote the Pauli operators acting on system $j$ and are defined in terms of
$X\coloneqq\begin{bmatrix}0&1\\ 1&0\end{bmatrix},\qquad Y\coloneqq\begin{bmatrix}0&-i\\ i&0\end{bmatrix},\qquad Z\coloneqq\begin{bmatrix}1&0\\ 0&-1\end{bmatrix}$. (8)
A parameterized thermal state of the thermodynamic system described by $H$ and $(Q_{1},\ldots,Q_{c})$, at temperature $T>0$, is given by
$\rho_{T}(\mu)\coloneqq\frac{\exp\!\left(-\frac{1}{T}\left(H-\mu\cdot Q\right)\right)}{Z_{T}(\mu)},$ (9)
where $\mu\coloneqq\left(\mu_{1},\ldots,\mu_{c}\right)\in\mathbb{R}^{c}$ is the chemical potential vector, we have used the shorthand
$\mu\cdot Q\coloneqq\sum_{i=1}^{c}\mu_{i}Q_{i},$ (10)
and the partition function $Z_{T}(\mu)$ is defined as
$Z_{T}(\mu)\coloneqq\operatorname{Tr}\!\left[\exp\!\left(-\frac{1}{T}\left(H-\mu\cdot Q\right)\right)\right].$ (11)
The chemical potential vector $\mu$ parameterizes the thermal state $\rho_{T}(\mu)$ and can be used for enabling $\rho_{T}(\mu)$ to satisfy constraints on expectation values of the charges in $Q$ (see, e.g., [16, Appendix B]). The state $\rho_{T}(\mu)$ is known in the quantum thermodynamics literature as a non-Abelian thermal state [22, 23, 24] and in the quantum machine learning literature as a quantum Boltzmann machine [25, 26, 27]. Let us note that the state $\rho_{T}(\mu)$ in (9) is well defined even if the charges are not conserved, but much of the focus in the quantum thermodynamics literature is on the case in which they are conserved (i.e., Eq. (4) is satisfied).
A low-temperature thermal state is an approximation of a Hamiltonian’s ground state, and vice versa, a Hamiltonian’s ground state is an approximation of a low-temperature thermal state. In Appendix A, we provide quantitative forms of this approximation in terms of three well known distinguishability measures for quantum states, including normalized trace distance, fidelity, quantum relative entropy, and quantum Rényi relative entropies.
All of the thermodynamic models that we deal with in this paper, including quantum Heisenberg models and stabilizer thermodynamic systems, satisfy the charge conservation constraint in (4). However, we should note that the developments in [16] do not depend on this constraint holding, nor does the interpretation of the LMPW algorithms given in Section III.
II.2 LMPW algorithms for constrained energy minimization
In this section, we review the problem of constrained energy minimization, as well as the LMPW classical and HQC algorithms for solving it [16].
II.2.1 Review of constrained energy minimization
Let $H$ be a Hamiltonian, and let $Q\coloneqq\left(Q_{1},\ldots,Q_{c}\right)$ be a tuple of non-commuting charges, where $c\in\mathbb{N}$ and $H,Q_{1},\ldots,Q_{c}$ are $d^{n}\times d^{n}$ Hermitian matrices, corresponding to observables for $n$ qudit systems. As mentioned previously, all of the developments in [16] do not require the charges in $Q$ to be conserved, and so we do not make such an assumption in this section.
Given $H$, $Q$, and a tuple $q\coloneqq\left(q_{1},\ldots,q_{c}\right)\in\mathbb{R}^{c}$ of constraint values, constrained energy minimization corresponds to the following optimization problem:
$E(\mathcal{Q},q)\coloneqq\min_{\rho\in\mathcal{D}_{d^{n}}}\left\{\operatorname{Tr}[H\rho]:\operatorname{Tr}[Q_{i}\rho]=q_{i}\ \forall i\in\left[c\right]\right\},$ (12)
where $\mathcal{D}_{d^{n}}$ denotes the set of $d^{n}\times d^{n}$ density matrices and $\mathcal{Q}\equiv\left(H,Q_{1},\ldots,Q_{c}\right)$. Clearly, not every tuple of constraints can be satisfied, in which case the value of $E(\mathcal{Q},q)$ is $\infty$. Going forward, in order to avoid this singularity, we therefore assume that the tuple of constraints, $\operatorname{Tr}[Q_{i}\rho]=q_{i}\ \forall i\in\left[c\right]$, is such that there exists at least one state $\rho$ satisfying them.
By introducing a temperature $T\geq 0$, one can approximate this problem by the following constrained free energy minimization problem [16, Section II-B]:
$F_{T}(\mathcal{Q},q)\coloneqq\\ \min_{\rho\in\mathcal{D}_{d^{n}}}\left\{\operatorname{Tr}[H\rho]-TS(\rho):\operatorname{Tr}[Q_{i}\rho]=q_{i}\ \forall i\in\left[c\right]\right\},$ (13)
where $S(\rho)\coloneqq-\operatorname{Tr}[\rho\ln\rho]$ is the von Neumann entropy. By invoking properties of von Neumann entropy, the following approximation holds [16, Lemma 4]:
$E(\mathcal{Q},q)\geq F_{T}(\mathcal{Q},q)\geq E(\mathcal{Q},q)-nT\ln d,$ (14)
so that one has $\varepsilon\geq 0$ additive error in the approximation by picking $T=\frac{\varepsilon}{n\ln d}$  [^1] Let us note that the dimension of the system we are considering here is $d^{n}$, instead of $d$ as considered previously in [16]. Observe that this choice implies that the temperature is low, i.e., proportional to the desired error $\varepsilon$ and inversely proportional to the number of qudits in the system.
By invoking Lagrangian duality and properties of quantum relative entropy, one can rewrite the constrained free energy minimization problem in terms of the following dual chemical potential maximization problem [16, Appendix B]:
$F_{T}(\mathcal{Q},q)=\sup_{\mu\in\mathbb{R}^{c}}f(\mu),$ (15)
where
$f(\mu)\coloneqq\mu\cdot q-T\ln Z_{T}(\mu)$ (16)
and the partition function $Z_{T}(\mu)$ is defined in (11). As part of the proof of (15), one deduces that parameterized thermal states of the form in (9) are optimal for the constrained free energy minimization problem in (13) (as mentioned before, these states are also known as non-Abelian thermal states or quantum Boltzmann machines). Additionally, the objective function $f(\mu)$ in (15) is concave in $\mu$ and the spectral norm of its Hessian is bounded from above by the following smoothness parameter [16, Lemma 9]:
$L\coloneqq\frac{2}{T}\sum_{i\in\left[c\right]}\left\|Q_{i}\right\|^{2}=\frac{2n\ln d}{\varepsilon}\sum_{i\in\left[c\right]}\left\|Q_{i}\right\|^{2},$ (17)
implying that standard optimization methods like gradient ascent and its variants [28] are guaranteed to converge to a globally optimal solution with a number of steps related to the smoothness parameter $L$ in (17).
The gradient $\nabla_{\mu}f(\mu)$ of the objective function $f(\mu)$ has the following entries [16, Eq. (18)]:
$\frac{\partial}{\partial\mu_{i}}f(\mu)=q_{i}-\operatorname{Tr}[Q_{i}\rho_{T}(\mu)].$ (18)
This form of the gradient allows for calculation by means of a classical algorithm for small system sizes or for estimation by means of quantum algorithm equipped with subroutines for thermal state preparation of $\rho_{T}(\mu)$ and measurement of each observable $Q_{i}$.
II.2.2 LMPW first-order classical algorithm
With the background from Section II.2.1 in hand, let us now recall the LMPW first-order classical algorithm for constrained energy minimization [16, Algorithm 1]:
Algorithm 1.
The algorithm proceeds according to the following steps:
  1. 1.
    Set $m\leftarrow 0$, and initialize $\mu^{m}\leftarrow\left(0,\ldots,0\right)$, the temperature $T\leftarrow\frac{\varepsilon}{n\ln d}$, the step size $\eta\in(0,L^{-1}]$, where $L$ is defined in (17), and the stopping parameter $\delta>0$.
  2. 2.
    Increment $m$, and set
    $\mu^{m}\leftarrow\mu^{m-1}+\eta\nabla_{\mu}f(\mu^{m-1}),$ (19)
    where the gradient $\nabla_{\mu}f(\mu^{m-1})$ is defined from (18).
  3. 3.
    Repeat Step 2 until $\left\|\nabla_{\mu}f(\mu^{m})\right\|\leq\delta$.
  4. 4.
    Output
    $\mu^{m}\cdot q+\operatorname{Tr}[\left(H-\mu^{m}\cdot Q\right)\rho_{T}(\mu^{m})]$ (20)
    as an approximation of $E(\mathcal{Q},q)$ in (12).
Recall from [16, Eq. (D2)] that the following equality holds:
$f(\mu)=\mu\cdot q+\operatorname{Tr}[\left(H-\mu\cdot Q\right)\rho_{T}(\mu)]-TS(\rho_{T}(\mu)),$ (21)
so that the output of Algorithm 1 in (20) is an approximation of $f(\mu^{m})$ that omits the last term $TS(\rho_{T}(\mu))$ in (21). As stated around [16, Eq. (21)], we could alternatively output $f(\mu^{m})$ as an approximation of the desired solution $E(\mathcal{Q},q)$, but the output in (20) is consistent with the output of the LMPW HQC algorithms. For our purposes here, this allows for a more direct comparison of the classical first-order algorithm above and the first-order HQC algorithm recalled in Section II.2.3.
We have phrased the LMPW classical algorithm slightly differently from how it was presented previously in [16, Algorithm 1], in order to simplify its presentation. The main difference is the introduction of the stopping parameter $\delta$ above, as a simple way of ending the algorithm after convergence has been achieved (i.e., when the norm of the gradient is small, further changes to $\mu$ are small and convergence near a globally optimal solution is guaranteed at this point by [77, Corollary 3.5]).
To be implemented on a classical computer, Algorithm 1 requires computation of matrix multiplications, traces, and matrix exponentials in order to evaluate the gradient in (18) and the algorithm’s output in (20), while adhering to the approach from (2) for encoding the numerical values of the thermal states. This is possible using standard tools like NumPy in Python [78], and it is what we used in our realizations of the classical algorithm, the results of which are reported in Section VI.
Let us also note that variations of gradient ascent can be employed instead of the standard approach presented in Algorithm 1. These include momentum [79], ADAM [80], and Nesterov accelerated gradient [81]. However, the preferred approach here is actually Nesterov accelerated gradient, because it converges most quickly among all of these methods and is in fact provably optimal whenever the objective function is smooth and concave [81], as it is in the constrained free-energy minimization problem.
Before moving on to the LMPW first-order HQC algorithm reviewed in Section II.2.3, let us note that the LMPW first-order classical algorithm can be considered a noiseless simulation of the first-order HQC algorithm. The main reason for this is that the classical algorithm numerically computes the expectation values $\operatorname{Tr}[H\rho_{T}(\mu)]$ and $\operatorname{Tr}[Q_{i}\rho_{T}(\mu)]$, for all $i\in\left[c\right]$, whereas the HQC algorithm estimates these expectation values through sampling on a quantum computer. As such, the classical algorithm can serve as a benchmark for the HQC algorithm, in addition to the benchmark provided by standard SDP solvers.
II.2.3 LMPW first-order hybrid quantum-classical algorithm
As proposed in [16], Algorithm 1 can be modified to a first-order HQC algorithm by replacing the gradient and output calculations in (19) and (20), respectively, with gradient and output estimations that are realized by a quantum computer. Estimating these quantities requires two capabilities: preparing the thermal state $\rho_{T}(\mu)$ of temperature $T$ and measuring the observables $H,Q_{1},\ldots,Q_{c}$. The former task is the bottleneck for the algorithm, given that measuring $H,Q_{1},\ldots,Q_{c}$ can be performed efficiently for observables of physical interest. As mentioned in Section I.3.1, here we focus on a basic thermal state preparation routine: direct preparation from the description of the density matrix from a quantum circuit that encodes its numerical values, while leaving it as an open direction to implement more advanced methods.
Algorithm 1 is then modified as follows, to become the LMPW first-order HQC algorithm for constrained energy minimization:
Algorithm 2.
The algorithm proceeds according to the following steps:
  1. 1.
    Set $m\leftarrow 0$, and initialize $\mu^{m}\leftarrow\left(0,\ldots,0\right)$, the temperature $T\leftarrow\frac{\varepsilon}{n\ln d}$, the step size $\eta\in(0,L^{-1}]$, where $L$ is defined in (17), and the stopping parameter $\delta>0$.
  2. 2.
    Increment $m$. For all $i\in\left[c\right]$, set $\widetilde{Q}_{i}$ as an estimate of $\operatorname{Tr}[Q_{i}\rho_{T}(\mu^{m-1})]$, and set
    $\mu_{i}^{m}\leftarrow\mu_{i}^{m-1}+\eta\left(q_{i}-\widetilde{Q}_{i}\right).$ (22)
  3. 3.
    Repeat Step 2 until $\sqrt{\sum_{i\in\left[c\right]}\left(q_{i}-\widetilde{Q}_{i}\right)^{2}}\leq\delta$.
  4. 4.
    Set $\widetilde{H}$ as an estimate of $\operatorname{Tr}[H\rho_{T}(\mu^{m})]$, and set $\widetilde{Q}_{i}$ as an estimate of $\operatorname{Tr}[Q_{i}\rho_{T}(\mu^{m})]$. Output
    $\mu^{m}\cdot q+\widetilde{H}-\sum_{i\in\left[c\right]}\mu_{i}^{m}\widetilde{Q}_{i}$ (23)
    as an approximation of $E(\mathcal{Q},q)$ in (12).
As presented above, Algorithm 2 is a simplification of the first-order HQC algorithm proposed in [16, Algorithm 2]. Since the gradient calculations are replaced by gradient estimates, the latter are random variables, and the algorithm thus employs stochastic gradient ascent in its search for a globally optimal solution (see, e.g., [77, Chapter 5] and [28, Section 6]). Since one of the goals of [16] was to prove that the algorithm converges to a global optimum, the presentation of the algorithm was more involved in order to provide a theoretical guarantee on its performance. However, in practice, we find in our simulations that the above simplification of the algorithm works well for converging near the global optimum.
The first-order HQC algorithm proposed in [16, Algorithm 2] focused on the case in which the Hamiltonian $H$ and the non-commuting charges $Q_{1},\ldots,Q_{c}$ can be written as a real linear combination of Pauli matrices. As described there, one can invoke a Monte Carlo sampling procedure in order to measure these observables. Such a procedure is useful when the coefficients in the linear combination of Pauli matrices differ markedly, because terms with less weight will be sampled less frequently, while terms with more weight will be sampled more frequently. However, in all of our example problems, the coefficients have equal weight, so that there is no benefit from performing such a Monte Carlo sampling procedure and we instead simply measure each term in the observables sufficiently many times in order to obtain reasonable estimates.
II.2.4 LMPW second-order classical algorithm
It is possible to modify Algorithm 1 to become a second-order classical algorithm, i.e., incorporating second-derivative information to aid the search [16, Section VI]. In this case, the Hessian of the objective function $f(\mu)$ has the following matrix elements [16, Eqs. (24)–(25)]:
$\frac{\partial^{2}}{\partial\mu_{i}\partial\mu_{j}}f(\mu)=-\frac{1}{T}\int_{0}^{1}ds\ \operatorname{Tr}[\rho_{T}(\mu)^{1-s}Q_{i}\rho_{T}(\mu)^{s}Q_{j}]\\ +\frac{1}{T}\operatorname{Tr}[Q_{i}\rho_{T}(\mu)]\operatorname{Tr}[Q_{j}\rho_{T}(\mu)].$ (24)
As noted in [16, Lemmas 6 and 8], the Hessian is negative semi-definite, implying that the objective function $f(\mu)$ in (16) is concave. Let us denote the Hessian matrix evaluated at $\nu\in\mathbb{R}^{c}$ by $\nabla_{\mu}^{2}f(\nu)$, so that it has matrix elements
$\left[\nabla_{\mu}^{2}f(\nu)\right]_{i,j}=-\frac{1}{T}\int_{0}^{1}ds\ \operatorname{Tr}[\rho_{T}(\nu)^{1-s}Q_{i}\rho_{T}(\nu)^{s}Q_{j}]\\ +\frac{1}{T}\operatorname{Tr}[Q_{i}\rho_{T}(\nu)]\operatorname{Tr}[Q_{j}\rho_{T}(\nu)].$ (25)
One can then use the explicit expression in (25) to calculate the Hessian at each iteration of Algorithm 1. This leads to the following modified classical algorithm for constrained energy minimization:
Algorithm 3.
The algorithm proceeds according to the following steps:
  1. 1.
    Set $m\leftarrow 0$, and initialize $\mu^{m}\leftarrow\left(0,\ldots,0\right)$, the temperature $T\leftarrow\frac{\varepsilon}{n\ln d}$, the step size $\eta>0$, and the stopping parameter $\delta>0$.
  2. 2.
    Increment $m$, and solve for $\Delta$ in
    $\nabla_{\mu}^{2}f(\mu^{m-1})\Delta=\nabla_{\mu}f(\mu^{m-1}),$ (26)
    where the gradient $\nabla_{\mu}f(\mu^{m-1})$ is defined from (18) and the Hessian matrix $\nabla_{\mu}^{2}f(\mu^{m-1})$ from (25). Set
    $\mu^{m}\leftarrow\mu^{m-1}-\eta\Delta.$ (27)
  3. 3.
    Repeat Step 2 until $\left\|\nabla_{\mu}f(\mu^{m})\right\|\leq\delta$.
  4. 4.
    Output
    $\mu^{m}\cdot q+\operatorname{Tr}[\left(H-\mu^{m}\cdot Q\right)\rho_{T}(\mu^{m})]$ (28)
    as an approximation of $E(\mathcal{Q},q)$ in (12).
Due to the Hessian being negative semi-definite, the minus sign in (27) effectively cancels out with the implicit minus sign from the Hessian $\nabla_{\mu}^{2}f(\mu^{m-1})$, so that each gradient update always proceeds in a positive direction and corresponds to an ascent procedure.
The algorithm can be sensitive to the step size $\eta$. The basic form of a Newton update simply sets $\eta=1$. However, such an update does not always lead to convergence, and instead we incorporate a backtracking approach based on the value of the norm of the gradient. If, at a certain iteration, the gradient norm increases, this indicates that the algorithm is not making progress, and we reduce the step size in a multiplicative way until the gradient norm decreases. At that point we continue with larger step sizes while making progress.
By incorporating second-derivative information, Algorithm 3 typically requires fewer iterations to converge than does Algorithm 1. However, each iteration of Algorithm 3 is computationally more expensive because it requires the calculation of the Hessian matrix.
Let us note that Algorithm 3 can be used as a direct benchmark for the performance of the second-order HQC algorithm detailed in Section II.2.5. As before, it can be considered a noiseless simulation of Algorithm 4, in the sense that it does not involve shot noise error that occurs from having to estimate the gradient and the Hessian of the objective function $f(\mu)$.
II.2.5 LMPW second-order hybrid quantum-classical algorithm
Finally, let us recall a second-order HQC algorithm for constrained energy minimization, as proposed in [16, Section VI]. A key observation, originally made in the context of a general-purpose natural gradient algorithm for quantum Boltzmann machines [82, Theorem 2], is that the Hessian matrix in (24) admits the following alternative expression [16, Lemma 6]:
$\frac{\partial^{2}}{\partial\mu_{i}\partial\mu_{j}}f(\mu)=-\frac{1}{2T}\operatorname{Tr}[\left\{\Phi_{\mu}(Q_{i}),Q_{j}\right\}\rho_{T}(\mu)]\\ +\frac{1}{T}\operatorname{Tr}[Q_{i}\rho_{T}(\mu)]\operatorname{Tr}[Q_{j}\rho_{T}(\mu)],$ (29)
where $\Phi_{\mu}$ is the following quantum channel [83, 84, 74, 85, 86, 82]:
$\Phi_{\mu}(\cdot)\coloneqq\int_{-\infty}^{\infty}dt\ p(t)\,e^{-i\left(H-\mu\cdot Q\right)t/T}(\cdot)e^{i\left(H-\mu\cdot Q\right)t/T},$ (30)
and $p(t)$ is the following high-peak tent probability density defined for $t\in\mathbb{R}$ [86, Eq. (9)]:
$p(t)\coloneqq\frac{2}{\pi}\ln\left|\coth(\pi t/2)\right|.$ (31)
The elements of the Hessian matrix in (29) break up into two terms (the first in the first line of (29) and the second in the second line of (29)). The second term can be estimated by measuring the observable $Q_{i}\otimes Q_{j}$ with respect to the state $\rho_{T}(\mu)\otimes\rho_{T}(\mu)$. The first term can be estimated by a combination of several subroutines, including classical random sampling from $p(t)$, Hamiltonian simulation [87, 88] according to $e^{-i\left(H-\mu\cdot Q\right)t/T}$, and the Hadamard test [89]. The precise procedure is detailed in [16, Figure 3 and Appendix H].
Let us now write down the explicit steps of the LMPW second-order HQC algorithm for constrained energy minimization:
Algorithm 4.
The algorithm proceeds according to the following steps:
  1. 1.
    Set $m\leftarrow 0$, and initialize $\mu^{m}\leftarrow\left(0,\ldots,0\right)$, the temperature $T\leftarrow\frac{\varepsilon}{n\ln d}$, the step size $\eta>0$, and the stopping parameter $\delta>0$.
  2. 2.
    Increment $m$. For all $i\in\left[c\right]$, set
    $\widetilde{\nabla}\leftarrow\left(q_{i}-\widetilde{Q}_{i}\right)_{i\in[c]}$ (32)
    as an estimate of the gradient, where $\widetilde{Q}_{i}$ is an estimate of $\operatorname{Tr}[Q_{i}\rho_{T}(\mu^{m-1})]$. Set $\widetilde{\nabla^{2}}$ as an estimate of the Hessian matrix in (29), by following [16, Algorithm 4]. Solve for $\Delta$ in
    $\widetilde{\nabla^{2}}\Delta=\widetilde{\nabla}$, (33)
    and set
    $\mu^{m}\leftarrow\mu^{m-1}-\eta\Delta,$ (34)
  3. 3.
    Repeat Step 2 until $\sqrt{\sum_{i\in\left[c\right]}\left(q_{i}-\widetilde{Q}_{i}\right)^{2}}\leq\delta$.
  4. 4.
    Set $\widetilde{H}$ as an estimate of $\operatorname{Tr}[H\rho_{T}(\mu^{m})]$, and set $\widetilde{Q}_{i}$ as an estimate of $\operatorname{Tr}[Q_{i}\rho_{T}(\mu^{m})]$ for all $i\in\left[c\right]$. Output
    $\mu^{m}\cdot q+\widetilde{H}-\sum_{i\in\left[c\right]}\mu_{i}^{m}\widetilde{Q}_{i}$ (35)
    as an approximation of $E(\mathcal{Q},q)$ in (12).
In addition to the backtracking approach mentioned after Algorithm 3, we also regularize the estimate of the Hessian matrix, by adding a scaled identity matrix to it, in order to ensure that it is negative semi-definite. By employing both backtracking and regularization, we find in our simulations that the algorithm converges in all cases. Without doing so, the estimate of the Hessian matrix need not be negative semi-definite, due to noise inherent in the sampling process, and the algorithm need not converge in such a case.
The implementation of the channel $\Phi_{\mu}$ can be simplified for estimating the Hessian when the charges are conserved (i.e., when (4) holds). To see this, observe that the term $\Phi_{\mu}(Q_{i})$ appears in (29), which can be written as follows:
$\Phi_{\mu}(Q_{i})$ $=\int_{-\infty}^{\infty}dt\ p(t)\,e^{-i\left(H-\mu\cdot Q\right)t/T}Q_{i}e^{i\left(H-\mu\cdot Q\right)t/T}$ (36)
$=\int_{-\infty}^{\infty}dt\ p(t)\,e^{i\mu\cdot Qt/T}Q_{i}e^{-i\mu\cdot Qt/T},$ (37)
where the second equality follows by employing the assumption in (4) and the fact that
$e^{A+B}=e^{A}e^{B}$ (38)
for Hermitian $A$ and $B$ satisfing $[A,B]=0$. If the charges are extensive in addition to being conserved (i.e., if (5) holds), the implementation of the channel $\Phi_{\mu}$ can be simplified even more. This is the case for the quantum Heisenberg models reviewed in Section II.3. In this case, it follows that
$\Phi_{\mu}(Q_{i})=\sum_{j=1}^{n}\int_{-\infty}^{\infty}dt\ p(t)\,e^{\frac{it}{T}\mu\cdot C^{(j)}}C_{i}^{(j)}e^{-\frac{it}{T}\mu\cdot C^{(j)}},$ (39)
where $C^{(j)}\equiv(C_{1}^{(j)},\ldots,C_{c}^{(j)})$. This simplification has the effect of reducing a multi-site Hamiltonian evolution, acting on all systems, to single-site evolutions, which act only on single systems, and can significantly reduce the circuit depth. See Appendix B for a proof of (39).
II.3 Quantum Heisenberg models
As mentioned in Section I.2, quantum Heisenberg models describe interacting spin systems on a lattice and are essential for understanding magnetic materials [38]. They are key examples considered for our simulations of the LMPW algorithms, the results of which we report in Section VI. A general form for a quantum Heisenberg model involves qubits placed at the vertices of a graph $G$ with vertex set $V$ and edge set $E$ [37]. The Hamiltonian is defined in terms of the graph $G$ and a list $\vec{J}\coloneqq(J_{i,j})_{\left\{i,j\right\}\in E}$ of coupling coefficients as follows:
$H(G,\vec{J})\coloneqq\\ \sum_{\left\{i,j\right\}\in E}J_{i,j}\left(X^{(i)}\otimes X^{(j)}+Y^{(i)}\otimes Y^{(j)}+Z^{(i)}\otimes Z^{(j)}\right).$ (40)
The total magnetizations in the $x$, $y$, and $z$ directions are defined as follows:
$X(G)$ $\coloneqq\sum_{i\in V}X^{(i)},$ (41)
$Y(G)$ $\coloneqq\sum_{i\in V}Y^{(i)},$ (42)
$Z(G)$ $\coloneqq\sum_{i\in V}Z^{(i)}.$ (43)
The following commutation relations hold for every graph $G$ and for every list $\vec{J}$ of coupling coefficients:
$\left[H(G,\vec{J}),X(G)\right]$ $=\left[H(G,\vec{J}),Y(G)\right]$ (44)
$=\left[H(G,\vec{J}),Z(G)\right]$ (45)
$=0.$ (46)
These are a direct consequence of the following commutation relations holding:
$\left[X\otimes X+Y\otimes Y+Z\otimes Z,X\otimes I+I\otimes X\right]$ $=0,$ (47)
$\left[X\otimes X+Y\otimes Y+Z\otimes Z,Y\otimes I+I\otimes Y\right]$ $=0,$ (48)
$\left[X\otimes X+Y\otimes Y+Z\otimes Z,Z\otimes I+I\otimes Z\right]$ $=0,$ (49)
which in turn follow directly from applying the equalities
$XY$ $=-YX=iZ,$ (50)
$YZ$ $=-ZY=iX,$ (51)
$ZX$ $=-XZ=iY.$ (52)
Clearly, the total magnetization operators $X(G)$, $Y(G)$, and $Z(G)$ do not form a commuting set.
As a consequence of the commutation relations in (44)–(46), it follows that every quantum Heisenberg model realizes a thermodynamic system with conserved, non-commuting charges, in which $H(G,\vec{J})$ is the Hamiltonian and $X(G)$, $Y(G)$, and $Z(G)$ are the conserved, non-commuting charges. Several examples of quantum Heisenberg models were closely studied in various prior works on quantum thermodynamics with conserved, non-commuting charges [23, 75, 24].
Special cases of the general model presented in (40)–(46) include the one- and two-dimensional quantum Heisenberg models with nearest-neighbor and next-to-nearest-neighbor interactions. In the case of a one-dimensional model, the qubits are arranged on a line with edges placed between nearest neighbors, for the nearest neighbor case, and edges additionally placed between next-to-nearest neighbors, for the next-to-nearest neighbor case. To simplify the model further, we set the coupling coefficients to be a fixed value $J$ between nearest neighbors, and for next-to-nearest neighbors, we set the coupling coefficients to be fixed at a constant value, say, $\lambda J$, where $\lambda\in[0,1]$. For a two-dimensional model, the main difference is that the qubits are arranged on a square lattice, and we take similar conventions, as mentioned above, for the nearest-neighbor and next-to-nearest neighbor cases. See Figure 1 for a depiction of these models.
Figure 1: Depiction of one- and two-dimensional Heisenberg models. (Top) Qubits of the one-dimensional model are arranged on a line with edges placed between nearest neighbors (black lines), and edges additionally placed between next-to-nearest neighbors (blue curves). (Bottom) Qubits of the two-dimensional model are arranged on a square lattice, with nearest neighbor and next-to-nearest neighbor interactions shown in black and blue lines, respectively.
II.4 Stabilizer codes for quantum error correction
Here we provide a brief review of stabilizer codes [44, 45] (see also [56]), which suffices for our purposes in this paper, and we point to various reviews for further exposition [39, 40, 41, 42, 43].
Let $\mathcal{P}^{n}$ denote the Pauli group for $n\in\mathbb{N}$ qubits. The Pauli group for a single qubit has the generating set $\left\langle X,Y,Z\right\rangle$, where $X$, $Y$, and $Z$ are defined in (8), and the Pauli group on $n$ qubits is generated by taking tensor products of the elements of this generating set.
Let $k\in\mathbb{N}$ be such that $k\leq n$. A stabilizer $\mathcal{S}$ is an Abelian subgroup of the Pauli group $\mathcal{P}^{n}$ not containing I-I and is generated by nkn-k commuting Pauli operators $S_{1},\ldots,S_{n-k}$:
$\mathcal{S}\coloneqq\left\langle S_{1},\ldots,S_{n-k}\right\rangle.$ (53)
By definition, the stabilizer operators $S_{1},\ldots,S_{n-k}$ in the generating set are independent, in the sense that it is not possible to realize any other one in the generating set by multiplying any two of them in it. The codespace $\mathcal{C}$ is the vector space stabilized by $\mathcal{S}$; i.e., it is equal to the simultaneous +1+1-eigenspace of the stabilizer operators in $\mathcal{S}$:
$\mathcal{C}\coloneqq\operatorname{span}\!\left\{|\psi\rangle\in\left(\mathbb{C}^{2}\right)^{\otimes n}:S|\psi\rangle=|\psi\rangle\ \ \forall S\in\mathcal{S}\right\}.$ (54)
The condition that $I\notin\mathcal{S}$ ensures that $\mathcal{S}$ has a simultaneous +1+1-eigenspace. The codespace $\mathcal{C}$ is the subspace of $\left(\mathbb{C}^{2}\right)^{\otimes n}$ into which quantum information is encoded. The projection onto the codespace is given by
$\Pi_{\mathcal{C}}$ $\coloneqq\frac{1}{2^{n-k}}\sum_{S\in\mathcal{S}}S=\prod_{i=1}^{n-k}\frac{I^{\otimes n}+S_{i}}{2},$ (55)
where the latter product is over $S_{1},\ldots,S_{n-k}$, i.e., those in the generating set for $\mathcal{S}$.
The set $\mathcal{L}$ of logical operators for the stabilizer $\mathcal{S}$ consists of the elements of the normalizer of $\mathcal{S}$:
$\mathcal{L}\coloneqq N(\mathcal{S})\coloneqq\left\{L\in\mathcal{P}^{n}:L\mathcal{S}=\mathcal{S}L\right\},$ (56)
where $L\mathcal{S}$ is the group realized by left multiplying all elements of $\mathcal{S}$ by $L$. For a stabilizer code, the normalizer $N(\mathcal{S})$ of $\mathcal{S}$ coincides with its centralizer [56, Section 3.2]:
$N(\mathcal{S})=C(\mathcal{S})\coloneqq\left\{L\in\mathcal{P}^{n}:LS=SL\ \ \forall S\in\mathcal{S}\right\}.$ (57)
The logical operators of $\mathcal{S}$ are generated by $2k$ operators that are labeled by $\overline{X}_{1},\overline{Z}_{1},\ldots,\overline{X}_{k},\overline{Z}_{k}$, with this notation being used because their commutation relations are the same as those of the standard Pauli operators $X_{1},Z_{1},\ldots,X_{k},Z_{k}$ acting on $k$ qubits:
$\left[\overline{X}_{i},\overline{X}_{j}\right]$ $=0\quad\forall i,j\in[k],$ (58)
$\left[\overline{Z}_{i},\overline{Z}_{j}\right]$ $=0\quad\forall i,j\in[k],$ (59)
$\left[\overline{X}_{i},\overline{Z}_{j}\right]$ $=0\quad\forall i,j\in[k]:i\neq j,$ (60)
$\left\{\overline{X}_{i},\overline{Z}_{i}\right\}$ $=0\quad\forall i\in[k].$ (61)
As such, the logical operators can be used to manipulate the $k$ logical qubits encoded in the stabilizer code. In correspondence with the equality $iXZ=Y$ that holds for the Pauli operators, we can also define the logical $\overline{Y}_{i}$ operator for each $i\in\left[k\right]$ as follows:
$\overline{Y}_{i}\coloneqq i\overline{X}_{i}\overline{Z}_{i}.$ (62)
A state $\rho$ of $k$ qubits is uniquely characterized by its Pauli coefficients $\left(r_{i_{1},\ldots,i_{k}}\right)_{i_{1},\ldots,i_{k}\in\{0,1,2,3\}}$:
$\rho=\frac{1}{2^{k}}\sum_{i_{1},\ldots,i_{k}\in\{0,1,2,3\}}r_{i_{1},\ldots,i_{k}}\sigma_{i_{1}}\otimes\cdots\otimes\sigma_{i_{k}},$ (63)
where we have introduced the notation $\sigma_{0}\equiv I$, $\sigma_{1}\equiv X$, $\sigma_{2}\equiv Y$, and $\sigma_{3}\equiv Z$. Similar to this, a state $\overline{\rho}$ of $k$ logical qubits encoded in the codespace $\mathcal{C}$ of a stabilizer code $\mathcal{S}$ is uniquely characterized by its logical Pauli coefficients $\left(r_{i_{1},\ldots,i_{k}}\right)_{i_{1},\ldots,i_{k}\in\{0,1,2,3\}}$:
$\overline{\rho}=\Pi_{\mathcal{C}}\left(\frac{1}{2^{k}}\sum_{i_{1},\ldots,i_{k}\in\{0,1,2,3\}}r_{i_{1},\ldots,i_{k}}\overline{\sigma}_{i_{1},1}\cdots\overline{\sigma}_{i_{k},k}\right)\Pi_{\mathcal{C}},$ (64)
where we have introduced the notation $\sigma_{0,i}\equiv\overline{I}_{i}=I^{\otimes n}$, $\sigma_{1,i}\equiv\overline{X}_{i}$, $\sigma_{2}\equiv\overline{Y}_{i}$, and $\sigma_{3}\equiv\overline{Z}_{i}$ for the logical operators of the code. For clarity, we include a short derivation of (64) in Appendix C.
III Design of Hamiltonians, ground states, and thermal states
In this section, we offer alternative interpretations and applications of the constrained energy minimization and constrained free energy minimization problems in (12) and (13), respectively. Namely, we show how applying the LMPW algorithms to these problems leads to the design of Hamiltonians with ground and thermal states satisfying fixed constraints on the expectation values of certain observables. In this way, the LMPW algorithms could be useful for the design of molecules and materials satisfying desired properties. Let us note that the design of ground states has been considered previously in the context of the density matrix renormalization group, in the special case that there is a single chemical potential that constrains a single conserved charge of the system [90, 91, 92].
To begin with, suppose that the Hamiltonian of a system of $n$ qudits is of the following form:
$H-\mu\cdot Q,$ (65)
where $H$ is an arbitrary $d^{n}\times d^{n}$ Hermitian matrix,
$Q\equiv\left(Q_{1},\ldots,Q_{c}\right)$ (66)
is a tuple of $d^{n}\times d^{n}$ Hermitian matrices, and $\mu\in\mathbb{R}^{c}$. Contrary to the setting of Section II.2.1, here we enforce no other constraints on $H$, $Q$, and $\mu$ (namely, we do not enforce the conservation constraint in (4)). Additionally, in distinction to the scenario put forth in Section II.2.1, now we no longer interpret $Q$ as being separate from the system Hamiltonian, but here instead we consider $Q$ to be a part of it. That is, we interpret $\mu$ as a vector of controllable parameters that one can vary in order to modify the Hamiltonian, such that it will satisfy fixed constraints on the expectations of the observables in $Q$.
As an example, the quantum Heisenberg model with an external magnetic field is as follows:
$H(G,\vec{J})-\mu_{x}X(G)-\mu_{y}Y(G)-\mu_{z}Z(G),$ (67)
where $H(G,\vec{J})$ is defined in (40), the operators $X(G)$, $Y(G)$, and $Z(G)$ are defined in (41)–(43), and $\left(\mu_{x},\mu_{y},\mu_{z}\right)\in\mathbb{R}^{3}$ is a vector of control parameters that determines the orientation and strength of an external magnetic field. One could then vary the control parameters $\mu_{x}$, $\mu_{y}$, and $\mu_{z}$ such that a ground or thermal state $\rho$ of (67) satisfies the following constraints on the total magnetizations:
$\operatorname{Tr}[X(G)\rho]=q_{x},\quad\operatorname{Tr}[Y(G)\rho]=q_{y},\quad\operatorname{Tr}[Z(G)\rho]=q_{z},$ (68)
where $\left(q_{x},q_{y},q_{z}\right)\in\mathbb{R}^{3}$ and under the assumption that there exists at least one state $\rho$ satisfying the constraints in (68).
III.1 Design of ground states
Let $q\equiv\left(q_{1},\ldots,q_{c}\right)\in\mathbb{R}^{c}$ be a vector of constraints. Suppose now that we would like to tune the control parameters $\mu$ of the Hamiltonian $H-\mu\cdot Q$, such that there is a state in its ground space satisfying the following constraints:
$\operatorname{Tr}[Q_{i}\rho]=q_{i}\qquad\forall i\in\left[c\right].$ (69)
We refer to this problem as “design of ground states.” As mentioned previously in Section II.2.1, not every tuple of constraints can be satisfied, in which case it is impossible to design a ground state (or any state for that matter) satisfying the constraints in (69). Thus, going forward from here, in order to avoid this singularity, we assume that there exists at least one state satisfying the tuple of constraints in (69).
Our claim is that this is possible by solving the following constrained energy minimization problem and its dual:
$\min_{\rho\in\mathcal{D}_{d^{n}}}\left\{\operatorname{Tr}[H\rho]:\operatorname{Tr}[Q_{i}\rho]=q_{i}\ \forall i\in\left[c\right]\right\}.$ (70)
To see this, let us suppose that there exists at least one state satisfying the constraints in (69), so that the minimum value in (70) is finite, and then determine the dual of the constrained energy minimization problem in (70) by applying the Lagrange multiplier method, as follows:
$\min_{\rho\in\mathcal{D}_{d^{n}}}\left\{\operatorname{Tr}[H\rho]:\operatorname{Tr}[Q_{i}\rho]=q_{i}\ \forall i\in\left[c\right]\right\}$
$=\min_{\rho\in\mathcal{D}_{d^{n}}}\left\{\operatorname{Tr}[H\rho]+\sup_{\mu\in\mathbb{R}^{c}}\sum_{i\in\left[c\right]}\mu_{i}\left(q_{i}-\operatorname{Tr}[Q_{i}\rho]\right)\right\}$ (71)
$=\min_{\rho\in\mathcal{D}_{d^{n}}}\sup_{\mu\in\mathbb{R}^{c}}\left\{\operatorname{Tr}[H\rho]+\sum_{i\in\left[c\right]}\mu_{i}\left(q_{i}-\operatorname{Tr}[Q_{i}\rho]\right)\right\}$ (72)
$=\min_{\rho\in\mathcal{D}_{d^{n}}}\sup_{\mu\in\mathbb{R}^{c}}\left\{\mu\cdot q+\operatorname{Tr}[\left(H-\mu\cdot Q\right)\rho]\right\}$ (73)
$=\sup_{\mu\in\mathbb{R}^{c}}\min_{\rho\in\mathcal{D}_{d^{n}}}\left\{\mu\cdot q+\operatorname{Tr}[\left(H-\mu\cdot Q\right)\rho]\right\}$ (74)
$=\sup_{\mu\in\mathbb{R}^{c}}\left\{\mu\cdot q+\lambda_{\min}(H-\mu\cdot Q)\right\}.$ (75)
The first equality in (71) follows by introducing $\mu$ as a vector of Lagrange multipliers to enforce the constraints in (69). The equality in (72) follows by moving the supremum to the outside, and the equality in (73) follows by simple algebra. The penultimate equality follows from the Sion minimax theorem for extended reals [93, Theorem 2.11]. The final equality follows by bringing in the minimum over all density matrices and recognizing that
$\min_{\rho\in\mathcal{D}_{d^{n}}}\operatorname{Tr}[\left(H-\mu\cdot Q\right)\rho]=\lambda_{\min}(H-\mu\cdot Q),$ (76)
where $\lambda_{\min}(A)$ denotes the minimum eigenvalue of a Hermitian operator $A$. In the theory of semi-definite programming, the equality between the first line and the last line above is known as strong duality [63, Section 5.9.1].
The development in (71)–(75) demonstrates that, in the case that the constrained energy minimization problem in (70) is feasible (i.e., there exists at least one state satisfying the constraints in (69)), then its minimum value is finite and, for the dual optimization problem in (75), there exists an optimal choice of $\mu$, which we denote by $\mu^{*}$. Furthermore, for the optimal choice $\mu^{*}$, there is a corresponding Hamiltonian, $H-\mu^{*}\cdot Q$, such that there exists a state in its ground space satisfying the constraints in (69). To see this, let $\rho^{*}$ be a state that is optimal for (70), and let $\mu^{*}$ be optimal for (75). Then, the complementary slackness conditions of semi-definite programming imply that the following equality holds:
$\left(H-\mu^{*}\cdot Q\right)\rho^{*}=\lambda_{\min}(H-\mu^{*}\cdot Q)\rho^{*}.$ (77)
See Appendix D for a derivation of (77). After taking a trace, Eq. (77) implies that $\rho^{*}$ is a state in the ground space of $H-\mu^{*}\cdot Q$.
In this way, by solving the constrained energy minimization problem in (70) and its dual in (75), we determine an optimal choice $\mu^{*}$ of the controllable parameters for the Hamiltonian $H-\mu^{*}\cdot Q$, such that there exists a state $\rho^{*}$ in the ground space of $H-\mu^{*}\cdot Q$ satisfying the constraints in (70). As such, by solving the primal and dual optimization problems, we can design Hamiltonians with a ground state satisfying desirable properties, i.e., the constraints in (69). The LMPW algorithms solve the constrained energy minimization problem approximately, by instead solving a free-energy minimization problem at low temperature, as reviewed in Section II.2, and preparing a low-temperature thermal state close to the ground state. After the LMPW HQC algorithm prepares this state, other observables could then be measured on it, in order to learn various properties of it.
Next, in Section III.2, we specifically discuss the free-energy minimization problem as a way of designing thermal states with desired properties.
III.2 Design of thermal states
The development from Section III.1 extends to the finite-temperature setting and the design of Hamiltonians having thermal states with desirable properties. We refer to this problem as “design of thermal states” and discuss it in more detail in this section.
As before, suppose that the controllable Hamiltonian is $H-\mu\cdot Q$, and the goal is find values of the control parameters in $\mu$ such that the temperature-$T$ thermal state of $H-\mu\cdot Q$ satisfies the constraints in (69). Recall that the temperature-$T$ thermal state of $H-\mu\cdot Q$ has the form given in (9).
Our claim is that this is possible by solving the dual of the following constrained free energy minimization problem:
$\min_{\rho\in\mathcal{D}_{d^{n}}}\left\{\operatorname{Tr}[H\rho]-TS(\rho):\operatorname{Tr}[Q_{i}\rho]=q_{i}\ \forall i\in\left[c\right]\right\}.$ (78)
To see this, let us revisit the derivation from [16, Appendix B], and before doing so, let us recall the quantum relative entropy $D(\omega\|\tau)$ of states $\omega$ and $\tau$ [94]:
$D(\omega\|\tau)$ $\coloneqq\operatorname{Tr}\!\left[\omega\left(\ln\omega-\ln\tau\right)\right]$ (79)
$=-S(\rho)-\operatorname{Tr}\!\left[\omega\ln\tau\right],$ (80)
with the property that $D(\omega\|\tau)\geq 0$ for all states $\omega$ and $\tau$ and $D(\omega\|\tau)=0$ if and only if $\omega=\tau$. As in the previous section, let us suppose that there exists at least one state satisfying the constraints in (69), so that the minimum value in (78) is finite. Now consider that
$\min_{\rho\in\mathcal{D}_{d^{n}}}\left\{\operatorname{Tr}[H\rho]-TS(\rho):\operatorname{Tr}[Q_{i}\rho]=q_{i}\ \forall i\in\left[c\right]\right\}$
$=\min_{\rho\in\mathcal{D}_{d^{n}}}\left\{\begin{array}[]{cc}\operatorname{Tr}[H\rho]-TS(\rho)+\\ \sup_{\mu\in\mathbb{R}^{c}}\sum_{i\in\left[c\right]}\mu_{i}\left(q_{i}-\operatorname{Tr}[Q_{i}\rho]\right)\end{array}\right\}$ (83)
$=\min_{\rho\in\mathcal{D}_{d^{n}}}\sup_{\mu\in\mathbb{R}^{c}}\left\{\begin{array}[]{cc}\operatorname{Tr}[H\rho]-TS(\rho)+\\ \sum_{i\in\left[c\right]}\mu_{i}\left(q_{i}-\operatorname{Tr}[Q_{i}\rho]\right)\end{array}\right\}$ (86)
$=\min_{\rho\in\mathcal{D}_{d^{n}}}\sup_{\mu\in\mathbb{R}^{c}}\left\{\mu\cdot q-TS(\rho)+\operatorname{Tr}[\left(H-\mu\cdot Q\right)\rho]\right\}$ (87)
$=\sup_{\mu\in\mathbb{R}^{c}}\min_{\rho\in\mathcal{D}_{d^{n}}}\left\{\mu\cdot q-TS(\rho)+\operatorname{Tr}[\left(H-\mu\cdot Q\right)\rho]\right\}$ (88)
$=\sup_{\mu\in\mathbb{R}^{c}}\min_{\rho\in\mathcal{D}_{d^{n}}}\left\{\mu\cdot q+TD(\rho\|\rho_{T}(\mu))-T\ln Z_{T}(\mu)\right\}$ (89)
$=\sup_{\mu\in\mathbb{R}^{c}}\left\{\mu\cdot q-T\ln Z_{T}(\mu)+T\min_{\rho\in\mathcal{D}_{d^{n}}}D(\rho\|\rho_{T}(\mu))\right\}$ (90)
$=\sup_{\mu\in\mathbb{R}^{c}}\left\{\mu\cdot q-T\ln Z_{T}(\mu)\right\}.$ (91)
The first four equalities follow from steps similar to the first few steps of (71)–(75). The equality in (89) follows from recognizing the quantum relative entropy, i.e.,
$-TS(\rho)+\operatorname{Tr}[\left(H-\mu\cdot Q\right)\rho]$
$=T\left(-S(\rho)-\operatorname{Tr}\!\left[\rho\left(-\frac{1}{T}\left(H-\mu\cdot Q\right)\right)\right]\right)$ (92)
$=T\left(-S(\rho)-\operatorname{Tr}\!\left[\rho\ln\exp\!\left(-\frac{1}{T}\left(H-\mu\cdot Q\right)\right)\right]\right)$ (93)
$=T\left(-S(\rho)-\operatorname{Tr}\!\left[\rho\ln\rho_{T}(\mu)\right]-\ln Z_{T}(\mu)\right)$ (94)
$=T\left[D(\rho\|\rho_{T}(\mu))-\ln Z_{T}(\mu)\right],$ (95)
where the last equality follows from (80). The final equality in (91) follows from the property of quantum relative entropy stated after (80).
The development in (83)–(91) demonstrates that, in the case that the constrained free energy minimization problem in (78) is feasible (i.e., there exists at least one state satisfying the constraints in (69)), then its minimum value is finite and, for the dual optimization problem in (91), there exists an optimal choice of $\mu$, which we denote by $\mu^{*}$. Furthermore, for the optimal choice $\mu^{*}$, there is a corresponding Hamiltonian, $H-\mu^{*}\cdot Q$, such that the thermal state $\rho_{T}(\mu^{*})$ satisfies the constraints in (69). To see this, observe that $(\mu^{*},\rho_{T}(\mu^{*}))$ is a saddle point of (83)–(91), such that all equalities are attained. Indeed, for the optimal choice $\mu^{*}$, the first-order optimality conditions for the objective function $f(\mu)\equiv\mu\cdot q-T\ln Z_{T}(\mu)$ in (91) guarantee that $\nabla_{\mu}f(\mu^{*})=0$, which implies by (18) that $q_{i}=\operatorname{Tr}[Q_{i}\rho_{T}(\mu^{*})]$ for all $i\in\left[c\right]$. Thus, the state $\rho_{T}(\mu^{*})$ satisfies all of the constraints. Using this condition, one then finds that
$\operatorname{Tr}[H\rho_{T}(\mu^{*})]-TS(\rho_{T}(\mu^{*}))=\mu^{*}\cdot q-T\ln Z_{T}(\mu^{*}).$ (96)
As such, due to the equalities in (83)–(91) holding, it follows that $\rho_{T}(\mu^{*})$ is a thermal state that is optimal for the constrained free energy minimization problem in (78).
In this way, by solving the constrained free energy minimization problem in (78) and its dual in (91), we determine the optimal choice $\mu^{*}$ of the controllable parameters for the Hamiltonian $H-\mu^{*}\cdot Q$, such that $\rho_{T}(\mu^{*})$ is a temperature-$T$ thermal state of $H-\mu^{*}\cdot Q$ satisfying the constraints in (70). As such, by solving this dual optimization problem, we can design Hamiltonians with thermal states satisfying desirable properties, i.e., the constraints in (69). The LMPW algorithms solve this problem to arbitrary accuracy and thus can be used for the design of thermal states. If one is exclusively interested in the design of thermal states, note that it is not necessary for the algorithms to output the final value of the free energy, but they instead can exclusively perform the gradient update steps in Algorithms 14 until the norm of the gradient is sufficiently small.
IV Stabilizer thermodynamic systems
Given the background in Sections II.1 and II.4, now we define a stabilizer thermodynamic system as follows:
Definition 5 (Stabilizer thermodynamic system).
Let $\mathcal{S}$ denote a stabilizer code that encodes $k$ logical qubits into $n$ physical qubits, has commuting stabilizer generators $S_{1},\ldots,S_{n-k}$, and has the set $\mathcal{L}$ of logical operators. We define a stabilizer thermodynamic system to have a Hamiltonian $H$ given by
$H\coloneqq-\sum_{i=1}^{n-k}\gamma_{i}S_{i},$ (97)
where $\gamma_{i}>0$, and conserved, non-commuting charges given by
$Q_{i}\coloneqq\sum_{j}\alpha_{i,j}L_{i,j},$ (98)
where $L_{i,j}\in\mathcal{L}$ and $\alpha_{i,j}\in\mathbb{R}$, for all $i\in\left[c\right]$ and $j$.
For the stabilizer Hamiltonian $H$ in (97), its ground space coincides with the codespace of $\mathcal{S}$, by construction [46]. Due to the fact that $\left[L_{i,j},S\right]=0$ for all $S\in\mathcal{S}$ (recall (56) and (57)), it follows that each $Q_{i}$ in (98) is a conserved charge, i.e., satisfying $\left[H,Q_{i}\right]=0$. Also, the tuple $\left(Q_{1},\ldots,Q_{c}\right)$ need not form a commuting set, as observed by recalling (58)–(61), in which case the charges are non-commuting in general.
It is clear from the definition in (98) that there is great freedom allowed in what is considered a conserved charge in a stabilizer thermodynamic system. As a first example, the conserved charges could be encoded total magnetizations, corresponding to the total magnetizations of the $k$ logical qubits, as follows:
$Q_{x}\coloneqq\sum_{i=1}^{k}\overline{X}_{i},\quad Q_{y}\coloneqq\sum_{i=1}^{k}\overline{Y}_{i},\quad Q_{z}\coloneqq\sum_{i=1}^{k}\overline{Z}_{i}.$ (99)
As another example, the conserved charges could be the following exhaustive list of all possible $4^{k}$ logical Pauli operators:
$Q_{i_{1},\ldots,i_{k}}\coloneqq\overline{\sigma}_{i_{1},1}\cdots\overline{\sigma}_{i_{k},k},$ (100)
where $i_{\ell}\in\{0,1,2,3\}$ and we have employed the notation introduced in (64). In this latter case, if all of the expectations $\left\langle Q_{i_{1},\ldots,i_{k}}\right\rangle_{\overline{\rho}}=r_{i_{1},\ldots,i_{k}}$ are fixed, where $\overline{\rho}$ is a state in the ground space of $H$ and $r_{i_{1},\ldots,i_{k}}\in\mathbb{R}$ for all $i_{1},\ldots,i_{k}\in\{0,1,2,3\}$, then the state of the $k$ logical qubits is completely specified after doing so (see (64)). This case leads to a particular way of using the LMPW HQC algorithms for encoding quantum information into the ground space of $H$, which we discuss in more detail in Section V.
Example 6.
As a simple example of a stabilizer thermodynamic system, let us consider the one-to-three qubit repetition code, which has a stabilizer generated by
$\left\langle Z_{1}Z_{2},Z_{2}Z_{3}\right\rangle,$ (101)
where $Z_{1}Z_{2}\equiv Z\otimes Z\otimes I$ and $Z_{2}Z_{3}\equiv I\otimes Z\otimes Z$, and logical operators generated by
$\left\langle X_{1}X_{2}X_{3},Z_{1}\right\rangle,$ (102)
where $X_{1}X_{2}X_{3}\equiv X\otimes X\otimes X$ and $Z_{1}\equiv Z\otimes I\otimes I$. In this case, the Hamiltonian of the corresponding stabilizer thermodynamic system is
$H=-Z_{1}Z_{2}-Z_{2}Z_{3},$ (103)
and we can take the conserved, non-commuting charges to be the encoded magnetizations of the logical qubit in the $x$, $y$, and $z$ directions:
$X_{\operatorname{tot}}$ $=X_{1}X_{2}X_{3},$ (104)
$Y_{\operatorname{tot}}$ $=iX_{1}X_{2}X_{3}Z_{1}=Y_{1}X_{2}X_{3},$ (105)
$Z_{\operatorname{tot}}$ $=Z_{1}.$ (106)
For the case of a single logical qubit, these operators coincide precisely with the $X$, $Y$, and $Z$ logical operators of the code, and constraints on them specify the state of the encoded qubit. The parameterized thermal state of this stabilizer thermodynamic system is as follows:
$\frac{e^{-\frac{1}{T}\left(-Z_{1}Z_{2}-Z_{2}Z_{3}-\mu_{x}X_{1}X_{2}X_{3}-\mu_{y}Y_{1}X_{2}X_{3}-\mu_{z}Z_{1}\right)}}{\operatorname{Tr}\!\left[e^{-\frac{1}{T}\left(-Z_{1}Z_{2}-Z_{2}Z_{3}-\mu_{x}X_{1}X_{2}X_{3}-\mu_{y}Y_{1}X_{2}X_{3}-\mu_{z}Z_{1}\right)}\right]},$ (107)
where $\left(\mu_{x},\mu_{y},\mu_{z}\right)\in\mathbb{R}^{3}$ is the chemical potential vector.
V Encoding quantum information into stabilizer thermodynamic systems
The LMPW first- and second-order HQC algorithms reviewed in Sections II.2.3 and II.2.5 can alternatively serve as methods for encoding quantum information into stabilizer thermodynamic systems. To see this, consider a stabilizer thermodynamic system corresponding to a stabilizer code encoding $k$ logical qubits into $n$ physical qubits, with a stabilizer $\mathcal{S}$ and a set $\mathcal{L}$ of logical operators. Suppose that we would like to encode a particular state of the $k$ logical qubits. Recall from (64) that such a state is specified by $4^{k}$ Pauli coefficients in the tuple $\left(r_{i_{1},\ldots,i_{k}}\right)_{i_{1},\ldots,i_{k}\in\{0,1,2,3\}}$. Then, after defining the Hamiltonian $H$ as in (97), with $\gamma_{i}=1$ for all $i\in[n-k]$, we set up a constrained energy minimization of the form in (12), as follows:
$\min_{\rho\in\mathcal{D}_{2^{n}}}\operatorname{Tr}[H\rho]$ (108)
subject to
$\operatorname{Tr}[\overline{\sigma}_{i_{1},1}\cdots\overline{\sigma}_{i_{k},k}\rho]$ $=r_{i_{1},\ldots,i_{k}}\>\forall i_{1},\ldots,i_{k}\in\{0,1,2,3\}.$
Evidently, the optimal solution to this problem is the value $\left(n-k\right)$, and the optimal state is $\overline{\rho}$, as given in (64). Indeed, Hamiltonian $H$ has a minimum eigenvalue of $\left(n-k\right)$, corresponding to a ground space that is $2^{k}$-fold degenerate. The constraints in (108) then uniquely specify the state of the logical qubits, according to (64).
By transforming this constrained energy minimization problem to a constrained free-energy minimization problem at temperature $T\geq 0$, as reviewed in Section II.2.1, the LMPW algorithms converge to an approximate solution of (108) for a sufficiently low temperature $T$. Furthermore, the HQC algorithms at the final step generate a parameterized thermal state of the following form:
$\frac{e^{-\frac{1}{T}\left(H-\sum_{i_{1},\ldots,i_{k}\in\{0,1,2,3\}}\mu_{i_{1},\ldots,i_{k}}\overline{\sigma}_{i_{1},1}\cdots\overline{\sigma}_{i_{k},k}\right)}}{\operatorname{Tr}\!\left[e^{-\frac{1}{T}\left(H-\sum_{i_{1},\ldots,i_{k}\in\{0,1,2,3\}}\mu_{i_{1},\ldots,i_{k}}\overline{\sigma}_{i_{1},1}\cdots\overline{\sigma}_{i_{k},k}\right)}\right]},$ (109)
where the chemical potentials in $\left(\mu_{i_{1},\ldots,i_{k}}\right)_{i_{1},\ldots,i_{k}\in\{0,1,2,3\}}$ ensure that the constraints in (108) are satisfied. Thus, in this way, the LMPW HQC algorithms can be viewed as methods for approximately encoding logical qubit states into stabilizer thermodynamic systems, whenever low-temperature thermal state preparation methods are experimentally accessible.
V.1 Warm-starting the encoding of a single qubit
The LMPW HQC algorithms start from the initial value $\mu=\left(0,\ldots,0\right)$, as seen in Algorithms 2 and 4. Such an initial starting point is used whenever there is no information about the solution to the problem at hand. However, in the case of encoding a single logical qubit into multiple physical qubits, we can provide a much better starting point for the algorithms, which speeds up the convergence significantly.
To see this, suppose that $\rho$ is a qubit state characterized by its Bloch vector $r\coloneqq\left(r_{x},r_{y},r_{z}\right)\in\mathbb{R}^{3}$:
$\rho=\frac{1}{2}\left(I+r_{x}X+r_{y}Y+r_{z}Z\right).$ (110)
If $\left\|r\right\|=1$, then $\rho$ is a pure state, and if $\left\|r\right\|\leq 1$, then $\rho$ is a mixed state. Whenever $\rho$ is a mixed state, it can alternatively be written in terms of a chemical potential vector $\mu\coloneqq\left(\mu_{x},\mu_{y},\mu_{z}\right)\in\mathbb{R}^{3}$ and an inverse temperature parameter $\beta\geq 0$ as follows:
$\rho=\frac{1}{Z_{\beta}(\mu)}\exp\!\left(-\beta\left(\mu_{x}X+\mu_{y}Y+\mu_{z}Z\right)\right),$ (111)
where
$Z_{\beta}(\mu)\coloneqq\operatorname{Tr}\!\left[\exp\!\left(-\beta\left(\mu_{x}X+\mu_{y}Y+\mu_{z}Z\right)\right)\right].$ (112)
The coordinates specified by $r$ are known as mixture coordinates, and those specified by $\mu$ and $\beta$ are known as exponential coordinates [95, Section 8.2]. The following theorem, proven in Appendix E, establishes a mapping from mixture coordinates to exponential coordinates for a single-qubit state:
Theorem 7.
Let $r$ be a Bloch vector corresponding to a qubit density matrix $\rho$, as written in (110), such that $\left\|r\right\|<1$. Then, by choosing
$\mu$ $=r,$ (113)
$\beta$ $=\frac{\operatorname{arctanh}\!\left(-\left\|r\right\|\right)}{\left\|r\right\|},$ (114)
the following equality holds:
$\frac{1}{2}\left(I+r_{x}X+r_{y}Y+r_{z}Z\right)=\\ \frac{1}{Z_{\beta}(\mu)}\exp\!\left(-\beta\left(\mu_{x}X+\mu_{y}Y+\mu_{z}Z\right)\right).$ (115)
Equipped with Theorem 7, we can initialize the LMPW HQC algorithms with the following state:
$\frac{e^{-\frac{1}{T}\left(H+\beta T\left(\mu_{x}\overline{X}+\mu_{y}\overline{Y}+\mu_{z}\overline{Z}\right)\right)}}{\operatorname{Tr}\!\left[e^{-\frac{1}{T}\left(H+\beta T\left(\mu_{x}\overline{X}+\mu_{y}\overline{Y}+\mu_{z}\overline{Z}\right)\right)}\right]},$ (116)
where $H$ is the stabilizer Hamiltonian for the code (as in (97), with $\gamma_{i}=1$ for all $i\in[n-k]$), $\overline{X}$, $\overline{Y}$, and $\overline{Z}$ are the logical operators for the encoded logical qubit, and $\mu$ and $\beta$ are chosen as in (113)–(114). The extra factor of $T$ multiplying $\beta$ in (116) ensures that the constraints in (108) are satisfied identically. In our simulations, we find that this initial starting point leads to immediate convergence of the LMPW HQC algorithms. In fact, the state in (116) is actually optimal, as a special case of the following more general theorem, which is proven in Appendix F:
Theorem 8.
Consider a stabilizer thermodynamic system with Hamiltonian
$H=-\sum_{i=1}^{n-k}S_{i},$ (117)
and logical operators $\overline{\sigma}_{i_{1},1}\cdots\overline{\sigma}_{i_{k},k}$ for all $i_{1},\ldots,i_{k}\in\left\{0,1,2,3\right\}$. Suppose that there is a mixed (non-pure) state $\rho$ of $k$ qubits satisfying the following constraints:
$\operatorname{Tr}\!\left[\left(\sigma_{i_{1}}\otimes\cdots\otimes\sigma_{i_{k}}\right)\rho\right]=r_{i_{1},\ldots,i_{k}}$ (118)
for all $i_{1},\ldots,i_{k}\in\left\{0,1,2,3\right\}$. (As indicated in (63), the constraints in (118) completely specify the $k$-qubit state.) Suppose that the coefficients in $\left(\mu_{i_{1},\ldots,i_{k}}\right)_{i_{1},\ldots,i_{k}\in\left\{0,1,2,3\right\}}$ are such that
$\rho=\frac{\exp\!\left(\sum_{i_{1},\ldots,i_{k}}\mu_{i_{1},\ldots,i_{k}}\sigma_{i_{1}}\otimes\cdots\otimes\sigma_{i_{k}}\right)}{\operatorname{Tr}\!\left[\exp\!\left(\sum_{i_{1},\ldots,i_{k}}\mu_{i_{1},\ldots,i_{k}}\sigma_{i_{1}}\otimes\cdots\otimes\sigma_{i_{k}}\right)\right]}.$ (119)
Then the following state
$\frac{\exp\!\left(-\frac{1}{T}\left(H-T\sum_{i_{1},\ldots,i_{k}}\mu_{i_{1},\ldots,i_{k}}\overline{\sigma}_{i_{1}}\cdots\overline{\sigma}_{i_{k}}\right)\right)}{\operatorname{Tr}\!\left[\exp\!\left(-\frac{1}{T}\left(H-T\sum_{i_{1},\ldots,i_{k}}\mu_{i_{1},\ldots,i_{k}}\overline{\sigma}_{i_{1}}\cdots\overline{\sigma}_{i_{k}}\right)\right)\right]}$ (120)
is the unique solution to the following constrained free energy minimization problem:
$\min_{\omega\in\mathcal{D}^{2^{n}}}\left\{\begin{array}[]{c}\operatorname{Tr}[H\omega]-TS(\omega):\\ \operatorname{Tr}\!\left[\left(\overline{\sigma}_{i_{1}}\cdots\overline{\sigma}_{i_{k}}\right)\omega\right]=r_{i_{1},\ldots,i_{k}}\end{array}\right\}.$ (121)
As expanded upon in Remark 18 of Appendix F, the intuition behind Theorem 8 is that, if the state of the logical qubits is completely specified by the constraints, then these degrees of freedom are not available for minimizing the free energy of the system, and one can only do so by initializing the ancilla qubits to temperature $T$ thermal states.
VI Simulation results
Figure 2: The figure depicts the logarithm of the error metric in (122) versus the logarithm of the number of iterations, for the task of constrained energy minimization for the two-dimensional, four-qubit quantum Heisenberg model with nearest- and next-to-nearest-neighbor interactions and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively. All of the algorithms converge, but the HQC algorithms, shown as the average over five independent runs with shaded regions denoting one standard deviation, require more iterations to converge due to sampling noise inherent in them.
In this section, we report the results of simulating the LMPW algorithms on several models of interest in quantum thermodynamics and the design of thermal states. Here, we focus on the one- and two-dimensional quantum Heisenberg models with a Hamiltonian having nearest- and next-to-nearest-neighbor interactions and the conserved, non-commuting charges set to the total magnetizations in the $x$, $y$, and $z$ directions (see Section II.3). We also report the results of simulations of a stabilizer thermodynamic system built from the perfect one-to-five qubit quantum error-correcting code.
In Appendix G, we report the results of other simulations, which include the one- and two-dimensional quantum Heisenberg models with nearest-neighbor interactions and similar settings as above for the charges. Therein, we also report simulation results for other stabilizer thermodynamic systems, including the one-to-three qubit repetition code and the two-to-four qubit quantum error-detecting code.
For all of our numerical experiments, we set the error metric for the constrained energy minimization problems to be as follows:
$\left|E^{\star}-\tilde{E}_{m}\right|+\left\|\tilde{\nabla}_{m}\right\|,$ (122)
in order to track the performance of the algorithms. In (122), $E^{\star}$ is the true minimum energy output from a standard SDP solver and $\tilde{E}_{m}$ and $\tilde{\nabla}_{m}$ are the estimates of the energy and the gradient, respectively, at the $m$th step of the algorithm. If $\left|E^{\star}-\tilde{E}_{m}\right|$ is small, then we are guaranteed that the algorithm has output a good approximation of the true energy at the $m$th step, and if $\left\|\tilde{\nabla}_{m}\right\|$ is small, then, according to (18), we are guaranteed that the algorithm has found a solution that satisfies the constraints in (12) approximately. In all of our figures, we plot the logarithm of the error metric in (122), in order to make it easier to visualize the performance of the algorithms. In addition to using this error metric for plotting purposes, we also used the gradient norm $\left\|\tilde{\nabla}_{m}\right\|$ as the stopping criterion to determine algorithm convergence, terminating the optimization once it fell below a specified threshold.
Figure 3: The figure depicts the logarithm of the error metric in (122) versus the number of iterations, for all of the LMPW algorithms (1st- and 2nd-order, and classical and HQC) for the task of constrained energy minimization for a stabilizer thermodynamic system formed from the perfect five-qubit error-correcting code. The constraints on the logical operators $\overline{X}$, $\overline{Y}$, and $\overline{Z}$ were set to $0.2$, 0, and $0.5$, respectively. All of the algorithms converge, but the HQC algorithms, shown as the average over five independent runs with shaded regions denoting one standard deviation, require more iterations to converge due to sampling noise inherent in them. For this simulation, we did not warm-start the algorithm according to the recipe from Section V.1, but we instead started with $\mu_{x}=1$, $\mu_{y}=1$, and $\mu_{z}=1$.
Our first set of simulations involves the two-dimensional quantum Heisenberg model with nearest- and next-to-nearest-neighbor interactions and with the charges set to the total magnetizations in the $x$, $y$, and $z$ directions. The goal is to solve the constrained energy minimization problem from Section II.2.1, with the expectation values on the total magnetizations in the $x$, $y$, and $z$ directions set to be $1$, 0, and $1$, respectively. Here we take the number of qubits to be fixed at four, arranged on a $2\times 2$ square lattice. Figure 2 compares the performance of all the LMPW algorithms, including the first- and second-order, classical and HQC algorithms, plotting the error metric in (122) versus the number of iterations. For the first- and second-order HQC algorithms, the average of five runs, along with one standard deviation shaded, is plotted. The figure demonstrates that all algorithms converge.
The classical and HQC second-order algorithms employ backtracking, as discussed after Algorithm 3. The HQC algorithms employ sampling, resulting in noisy estimates of the gradient and, for the second-order algorithm, the Hessian. We used approximately $10^{4}$ samples per iteration to estimate expectation values of observables in both the first-order and second-order HQC algorithms, and $10^{7}$ samples per iteration were employed to estimate the Hessian for the second-order HQC algorithm. To reduce the computational cost of Hessian estimation, we parallelized the simulation across multiple CPU cores using process-level parallelism. Each process evaluates a subset of Hadamard-test circuits with a dedicated, single-threaded simulator to avoid resource contention. The circuits are pre-transpiled with placeholders for the time-evolution unitaries, which are dynamically inserted at runtime using cached matrix exponentials. This strategy significantly accelerates the Hessian estimation step while keeping memory usage and inter-process contention low. As shown in the figure, the second-order HQC exhibits a higher variance across independent runs, primarily due to the statistical and numerical challenges involved in estimating and inverting the Hessian. In particular, the inverse Hessian estimate is significantly more sensitive to noise and often suffers from numerical instabilities, which can amplify fluctuations in the update direction. To mitigate the computational cost of Hessian estimation, we adopted a looser convergence threshold for the gradient norm in the second-order HQC algorithm. This explains the comparatively higher error metric reached by the second-order HQC algorithm: unlike the other algorithms, whose faster runtime permitted more iterations and finer convergence, the second-order HQC was terminated earlier to maintain computational feasibility.
The figure clearly illustrates that incorporating second-derivative information, as done in the second-order algorithms, improves performance, i.e., requiring much fewer iterations to reach convergence compared to the first-order methods. However, as stated previously, the second-order algorithms require more computations per iteration than the first-order algorithms, and thus a trade-off is present.
Our next set of simulations reported here involve a stabilizer thermodynamic system constructed from the perfect one-to-five qubit code. Specifically, recall that a generating set for the code’s stabilizer consists of the following operators [56, Section 8.2]:
$S_{1}$ $\coloneqq X\otimes Z\otimes Z\otimes X\otimes I,$ (123)
$S_{2}$ $\coloneqq I\otimes X\otimes Z\otimes Z\otimes X,$ (124)
$S_{3}$ $\coloneqq X\otimes I\otimes X\otimes Z\otimes Z,$ (125)
$S_{4}$ $\coloneqq Z\otimes X\otimes I\otimes X\otimes Z,$ (126)
and the code’s logical operators are as follows:
$\overline{X}$ $\coloneqq X\otimes X\otimes X\otimes X\otimes X,$ (127)
$\overline{Y}$ $\coloneqq Y\otimes Y\otimes Y\otimes Y\otimes Y,$ (128)
$\overline{Z}$ $\coloneqq Z\otimes Z\otimes Z\otimes Z\otimes Z.$ (129)
As such, the Hamiltonian for the stabilizer thermodynamic system is as follows:
$H\coloneqq-\left(S_{1}+S_{2}+S_{3}+S_{4}\right),$ (130)
and the conserved, non-commuting charges are set to $\overline{X}$, $\overline{Y}$, and $\overline{Z}$. In this case, the thermal state of the system at temperature $T$ is given by
$\frac{e^{-\frac{1}{T}\left(H-\mu_{x}\overline{X}-\mu_{y}\overline{Y}-\mu_{z}\overline{Z}\right)}}{\operatorname{Tr}[e^{-\frac{1}{T}\left(H-\mu_{x}\overline{X}-\mu_{y}\overline{Y}-\mu_{z}\overline{Z}\right)}]},$ (131)
where $\left(\mu_{x},\mu_{y},\mu_{z}\right)\in\mathbb{R}^{3}$ is the chemical potential vector. According to the discussion after (108), the ground energy of this stabilizer thermodynamic system is equal to (nk)=4-(n-k)=-4, because $n=5$ and $k=1$ in this case.
Figure 4: The figure depicts the logarithm of the error metric in (122) versus the number of iterations, for the task of constrained energy minimization for the one-dimensional quantum Heisenberg model with nearest- and next-to-nearest-neighbor interactions and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively. The LMPW classical algorithm converges in all cases (3, 4, 5, and 6 qubits), with larger systems taking more iterations to converge.
Figure 3 plots the performance of the LMPW algorithms for the above stabilizer thermodynamic system. Again the goal is to solve the constrained energy minimization problem from Section II.2.1. However, as mentioned in Section V, the LMPW HQC algorithms in this context can alternatively be understood as methods for encoding quantum information into an error-correcting code. Indeed, by setting the expectation values of the logical operators to be
$\operatorname{Tr}[\overline{X}\rho]$ $=r_{x},$ (132)
$\operatorname{Tr}[\overline{Y}\rho]$ $=r_{y},$ (133)
$\operatorname{Tr}[\overline{Z}\rho]$ $=r_{z},$ (134)
where $\left(r_{x},r_{y},r_{z}\right)\in\mathbb{R}^{3}$ is a vector in the Bloch ball, the LMPW HQC algorithms encode a logical qubit with this Bloch sphere representation. In our simulations, we choose
$r_{x}=0.2,\quad r_{y}=0,\quad r_{z}=0.5.$ (135)
The plot in Figure 3 is qualitatively similar to that from Figure 2, comparing the error metric in (122) versus number of iterations for all of the algorithms. The details of the implementation are similar to those of the previous example, and so we do not repeat them here. Again we find that the second-order algorithms require fewer iterations than do the first-order algorithms, however with the same trade-off mentioned previously.
Figure 5: The figure depicts the average of the logarithm of the error metric in (122) over five independent runs, plotted against the number of iterations, with shaded regions indicating one standard deviation. The task is constrained energy minimization for the one-dimensional quantum Heisenberg model with nearest- and next-to-nearest-neighbor interactions, and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively. The LMPW HQC algorithm converges in all cases (3, 4, 5, and 6 qubits), with larger systems taking more iterations to converge.
Our final set of simulations involves the one-dimensional quantum Heisenberg model with nearest- and next-to-nearest-neighbor interactions and with the charges set to the total magnetizations in the $x$, $y$, and $z$ directions. The goal is to solve the constrained energy minimization problem reviewed in Section II.2.1, with the expectation values on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively. Figure 4 demonstrates the performance of the first-order LMPW classical algorithm as the number of qubits in the model increases (specifically, the figure plots results for three, four, five, and six qubits). The figure plots the error metric in (122) versus the number of iterations, demonstrating that every case converges to the true value as the number of iterations increases.
Figure 5 demonstrates the performance of the first-order HQC algorithm for the same setup. The algorithm converges in all cases, but requires more iterations than the classical algorithm and exhibits noticeable noise in proximity to the optimal value. As expected, this behavior arises from the sampling noise inherent in the HQC algorithm. To mitigate the increased computational cost and instability associated with this noise, we employed a looser convergence threshold on the gradient norm, which also accounts for the slightly higher final error metric reached by the HQC algorithm compared to its classical counterpart. The number of shots used per iteration was set according to [16, Eq. (F17)], with $\varepsilon=10^{-4}$ and $\delta=10^{-2}$. In this case, the norm of the Pauli coefficients of the total magnetizations is $n$, where $n$ is the number of qubits.
VI.1 Comparison of Standard Gradient Ascent and Nesterov’s Accelerated Method
Figure 6: The figure depicts the logarithm of the error metric in (122) versus the number of iterations for the LMPW classical first-order algorithm, comparing the standard gradient ascent method with Nesterov’s accelerated version. The task is constrained energy minimization for the two-dimensional, four-qubit quantum Heisenberg model with nearest- and next-to-nearest-neighbor interactions and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively.
Figure 7: The figure depicts the average of the logarithm of the error metric in (122) over five independent runs, plotted against the number of iterations, with shaded regions indicating one standard deviation, for the LMPW HQC first-order algorithm, comparing the standard gradient ascent method with Nesterov’s accelerated version. The task is constrained energy minimization for the two-dimensional, four-qubit quantum Heisenberg model with nearest- and next-to-nearest-neighbor interactions and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively.
Throughout this work, we employed the standard gradient ascent algorithm for both the classical and HQC implementations of the first-order optimization procedure. However, several well-known modifications of the standard gradient ascent method can lead to improved convergence performance, especially in concave settings. One such modification is Nesterov’s accelerated gradient method [81], which introduces a momentum term to anticipate the direction of the next update, resulting in improved convergence speed.
While the standard gradient ascent update rule, given in (19), directly updates the parameters based on the gradient evaluated at the current iterate, Nesterov’s accelerated gradient ascent maintains an auxiliary sequence $v_{m}$, which incorporates a momentum term, and performs the update as follows:
$v_{m+1}$ $=\tau v_{m}+\eta\nabla f(\mu_{m}+\tau v_{m}),$ (136)
$\mu_{m+1}$ $=\mu_{m}+v_{m+1},$ (137)
where $\tau\in[0,1)$ is the momentum parameter and $\eta>0$ is the learning rate. The key idea is to compute the gradient not at the current iterate $\mu_{m}$, but at the projected position $\mu_{m}+\tau v_{m}$, effectively introducing a look-ahead mechanism that often leads to faster convergence.
Figure 6 illustrates the performance of the standard and Nesterov-accelerated gradient ascent algorithms for the classical first-order optimization of the two-dimensional four-qubit Heisenberg model with nearest- and next-to-nearest-neighbor interactions. While both methods converge to the same solution, the Nesterov-accelerated method does so in significantly fewer iterations. Figure 7 presents the corresponding results for the HQC first-order optimization applied to the same task, again demonstrating that the Nesterov-accelerated method converges in significantly fewer iterations than the standard gradient ascent method.
VII Conclusion
In this paper, we put forward several conceptual contributions, including framing constrained energy minimization as a method for designing ground states of Hamiltonians and constrained free energy minimization as a method for designing thermal states of Hamiltonians. We argued how the LMPW HQC algorithms can be used for designing these Hamiltonians and how this application has implications for designing molecules and materials with desired properties. We also introduced stabilizer thermodynamic systems as thermodynamic systems for which the Hamiltonian is constructed from the stabilizer operators of a stabilizer code and the conserved, non-commuting charges are constructed from the code’s logical operators. Then we demonstrated how the LMPW HQC algorithms can be understood as alternative methods for encoding quantum information into quantum error-correcting codes, which could be a useful approach whenever low-temperature thermal state preparation methods are experimentally accessible.
The numerical contribution of our paper involved simulating the LMPW algorithms for several key thermodynamic systems of interest, including the one- and two-dimensional quantum Heisenberg models with nearest-neighbor and next-to-nearest-neighbor interactions and several stabilizer thermodynamic systems (the one-to-three qubit repetition code, the two-to-four qubit error-detecting code, and the perfect one-to-five qubit code). To test the performance of the LMPW algorithms on all of these examples, we employed standard SDP solvers to determine the true values of the constrained energy minimization problems. The LMPW classical algorithms can be understood as establishing alternative benchmarks for the HQC algorithms: the former algorithms act as noiseless simulators of the HQC algorithms, which focus exclusively on converging when measurement shot noise is not present, while our simulations of the latter incorporate measurement shot noise.
Going forward from here, there are several avenues to pursue for future research. First, we think it would be worthwhile to employ matrix product states for simulating the thermal states needed in the LMPW classical algorithms, as used in prior work [96, 67, 97]. Very likely such an approach would lead an advanced classical algorithm and serve several purposes: first, it could be a useful algorithm on its own for understanding quantum thermodynamic systems; second, it could be helpful for solving or approximating the design of ground and thermal states of controllable Hamiltonians; third, it could be useful as a benchmark to set against the LMPW HQC algorithms. Regarding the latter point, any claimed quantum speedup for the LMPW HQC algorithms would have to outperform any possible classical algorithm for this task, and it is likely that approaches involving matrix product states would lead to the fastest possible classical algorithms.
We also left it open to test the performance of the LMPW HQC algorithms on currently available quantum computers. In doing so, it would be necessary to incorporate the techniques of quantum error correction and mitigation in order to limit the effects of noise, and one would also need to incorporate various advanced methods for thermal state preparation [29, 30, 31, 32, 33, 34, 35, 36] in order to have efficient approaches for this necessary step. As discussed in [16], in order to have an efficient HQC algorithm, it would be necessary to find a quantum thermodynamic system (Hamiltonian and conserved charges) for which one can efficiently prepare low-temperature thermal states.
Acknowledgements.
We are grateful to Nicole Yunger Halpern for a helpful discussion on quantum thermodynamics in the presence of conserved, non-commuting charges. We also thank Zoë Holmes, Peter McMahon, Karan Mehta, Erich Mueller, and Darren Pereira for helpful discussions. MM acknowledges support of the NCCR MARVEL, a National Centre of Competence in Research, funded by the Swiss National Science Foundation (grant number 205602). MC and JK acknowledge support from the Cornell Engineering Learning Initiative program. IL, ML, and KW are grateful to the Cornell School of Electrical and Computer Engineering for hospitality during their summer internships. NL acknowledges funding from the Science and Technology Commission of Shanghai Municipality (STCSM) grant no. 24LZ1401200 (21JC1402900), NSFC grants no. 12471411 and no. 12341104, the Shanghai Jiao Tong University 2030 Initiative, and the Fundamental Research Funds for the Central Universities. MMW acknowledges support from the National Science Foundation under grant no. 2329662 and from the Cornell School of Electrical and Computer Engineering.
Author Contributions
Author Contributions: The following describes the different contributions of the authors of this work, using roles defined by the CRediT (Contributor Roles Taxonomy) project [98]:
MM: Data curation, Investigation, Resources, Software, Supervision, Validation, Visualization, Writing - review & editing.
MC: Data curation, Resources, Software, Validation, Visualization, Writing - review & editing.
JK: Data curation, Resources, Software, Validation, Visualization, Writing - review & editing.
NL: Conceptualization, Methodology, Writing - review & editing.
IL: Data curation, Resources, Software, Validation, Visualization, Writing - review & editing.
ML: Data curation, Resources, Software, Validation, Visualization.
SR: Software, Supervision, Validation.
KW: Data curation, Resources, Software, Validation, Visualization, Writing - review & editing.
MMW: Conceptualization, Formal Analysis, Funding acquisition, Investigation, Methodology, Project Administration, Supervision, Validation, Writing - original draft, Writing - review & editing.
Appendix A Closeness of low-temperature thermal states and ground states of Hamiltonians
How well does the thermal state of a Hamiltonian approximate its ground state? Alternatively, in the case of a degenerate Hamiltonian, how well does the thermal state of a Hamiltonian approximate the maximally mixed state of its ground space? It is well known that the approximation becomes better as $\beta\to\infty$, where $\beta\coloneqq\frac{1}{T}$ is the inverse temperature.
In this appendix, we answer this question quantitatively in terms of several standard distinguishability measures for quantum states: the normalized trace distance (Proposition 9), the fidelity (Proposition 11), the quantum relative entropy, and quantum Rényi relative entropies (Proposition 13). Recall that these measures are respectively defined for states $\rho$ and $\sigma$ and $\alpha\in(0,1)\cup(1,\infty)$ as follows:
$\frac{1}{2}\left\|\rho-\sigma\right\|_{1}$ , (138)
$F(\rho,\sigma)$ $\coloneqq\left\|\sqrt{\rho}\sqrt{\sigma}\right\|_{1}^{2},$ (139)
$D(\rho\|\sigma)$ $\coloneqq\operatorname{Tr}[\rho\left(\ln\rho-\ln\sigma\right)],$ (140)
$D_{\alpha}(\rho\|\sigma)$ $\coloneqq\frac{1}{\alpha-1}\ln\operatorname{Tr}[\rho^{\alpha}\sigma^{1-\alpha}],$ (141)
where $\left\|A\right\|_{1}\coloneqq\operatorname{Tr}\!\left[\sqrt{A^{{\dagger}}A}\right]$. The Rényi relative entropy in (141) is the Petz–Rényi relative entropy [99], and we note that there are other quantum Rényi relative entropies, including the sandwiched Rényi relative entropy [100, 101] and the geometric Rényi relative entropy [102, 103], defined for $\alpha\in(0,1)\cup(1,\infty)$ as follows:
$\widetilde{D}_{\alpha}(\rho\|\sigma)$ $\coloneqq\frac{1}{\alpha-1}\ln\operatorname{Tr}\!\left[\left(\sigma^{\frac{1-\alpha}{2\alpha}}\rho\sigma^{\frac{1-\alpha}{2\alpha}}\right)^{\alpha}\right],$ (142)
$\widehat{D}_{\alpha}(\rho\|\sigma)$ $\coloneqq\frac{1}{\alpha-1}\ln\operatorname{Tr}\!\left[\sigma\left(\sigma^{-\frac{1}{2}}\rho\sigma^{-\frac{1}{2}}\right)^{\alpha}\right].$ (143)
For our case of interest here, the states being considered commute, i.e., $\left[\rho,\sigma\right]=0$, in which case the following equalities hold for $\alpha\in(0,1)\cup(1,\infty)$:
$D_{\alpha}(\rho\|\sigma)=\widetilde{D}_{\alpha}(\rho\|\sigma)=\widehat{D}_{\alpha}(\rho\|\sigma).$ (144)
Let $H$ be a Hamiltonian for a $d$-dimensional quantum system with the following spectral decomposition:
$H=\sum_{i=1}^{M}\lambda_{i}\Pi_{i},$ (145)
where we assume that the eigenvalues are ordered, so that $\lambda_{M}\geq\cdots\geq\lambda_{1}$, and $\left\{\Pi_{i}\right\}_{i=1}^{M}$ is the set of spectral projections, satisfying $\Pi_{i}\Pi_{j}=\delta_{i,j}\Pi_{i}$ for all $i,j\in\left[M\right]$ and $\sum_{i=1}^{M}\Pi_{i}=I$.
Proposition 9.
For a $d$-dimensional Hamiltonian $H$ as in (145) and $\beta\geq 0$, the following hold:
$\frac{1}{2}\left\|\frac{e^{-\beta H}}{\operatorname{Tr}[e^{-\beta H}]}-\frac{\Pi_{1}}{\operatorname{Tr}[\Pi_{1}]}\right\|_{1}$
$=\frac{1}{1+\frac{d_{G}}{\sum_{i=2}^{M}e^{-\beta\left(\lambda_{i}-\lambda_{1}\right)}\operatorname{Tr}[\Pi_{i}]}}.$ (146)
$\leq\frac{1}{1+e^{\beta\Delta}\frac{d_{G}}{d-d_{G}}},$ (147)
where $\Delta\coloneqq\lambda_{2}-\lambda_{1}$ is the spectral gap and $d_{G}\coloneqq\operatorname{Tr}[\Pi_{1}]$ is the dimension of the ground space of $H$.
Proof.
Consider that
$\frac{e^{-\beta H}}{\operatorname{Tr}[e^{-\beta H}]}=\frac{\sum_{i=1}^{M}e^{-\beta\lambda_{i}}\Pi_{i}}{\sum_{i=1}^{M}e^{-\beta\lambda_{i}}\operatorname{Tr}[\Pi_{i}]}.$ (148)
Now consider that
$\left\|\frac{\sum_{i=1}^{M}e^{-\beta\lambda_{i}}\Pi_{i}}{\sum_{i=1}^{M}e^{-\beta\lambda_{i}}\operatorname{Tr}[\Pi_{i}]}-\frac{\Pi_{1}}{\operatorname{Tr}[\Pi_{1}]}\right\|_{1}$
$=\left\|\frac{\sum_{i=2}^{M}e^{-\beta\lambda_{i}}\Pi_{i}}{\sum_{i=1}^{M}e^{-\beta\lambda_{i}}\operatorname{Tr}[\Pi_{i}]}+\frac{e^{-\beta\lambda_{1}}\Pi_{1}}{\sum_{i=1}^{M}e^{-\beta\lambda_{i}}\operatorname{Tr}[\Pi_{i}]}-\frac{\Pi_{1}}{\operatorname{Tr}[\Pi_{1}]}\right\|_{1}$ (149)
$=\left\|\begin{array}[]{cc}\frac{\sum_{i=2}^{M}e^{-\beta\lambda_{i}}\Pi_{i}}{\sum_{i=1}^{M}e^{-\beta\lambda_{i}}\operatorname{Tr}[\Pi_{i}]}+\\ \left(\frac{e^{-\beta\lambda_{1}}}{\sum_{i=1}^{M}e^{-\beta\lambda_{i}}\operatorname{Tr}[\Pi_{i}]}-\frac{1}{\operatorname{Tr}[\Pi_{1}]}\right)\Pi_{1}\end{array}\right\|_{1}$ (152)
$=\left\|\frac{\sum_{i=2}^{M}e^{-\beta\lambda_{i}}\Pi_{i}}{\sum_{i=1}^{M}e^{-\beta\lambda_{i}}\operatorname{Tr}[\Pi_{i}]}\right\|_{1}$
$\qquad+\left|\frac{e^{-\beta\lambda_{1}}}{\sum_{i=1}^{M}e^{-\beta\lambda_{i}}\operatorname{Tr}[\Pi_{i}]}-\frac{1}{\operatorname{Tr}[\Pi_{1}]}\right|\left\|\Pi_{1}\right\|_{1}$ (153)
$=\frac{\sum_{i=2}^{M}e^{-\beta\lambda_{i}}\left\|\Pi_{i}\right\|_{1}}{\sum_{i=1}^{M}e^{-\beta\lambda_{i}}\operatorname{Tr}[\Pi_{i}]}$
$\qquad+\left|\frac{e^{-\beta\lambda_{1}}}{\sum_{i=1}^{M}e^{-\beta\lambda_{i}}\operatorname{Tr}[\Pi_{i}]}-\frac{1}{\operatorname{Tr}[\Pi_{1}]}\right|\operatorname{Tr}[\Pi_{1}]$ (154)
$=\frac{\sum_{i=2}^{M}e^{-\beta\lambda_{i}}\operatorname{Tr}[\Pi_{i}]}{\sum_{i=1}^{M}e^{-\beta\lambda_{i}}\operatorname{Tr}[\Pi_{i}]}+\left|\frac{e^{-\beta\lambda_{1}}\operatorname{Tr}[\Pi_{1}]}{\sum_{i=1}^{M}e^{-\beta\lambda_{i}}\operatorname{Tr}[\Pi_{i}]}-1\right|$ (155)
$=\frac{\sum_{i=2}^{M}e^{-\beta\lambda_{i}}\operatorname{Tr}[\Pi_{i}]}{\sum_{i=2}^{M}e^{-\beta\lambda_{i}}\operatorname{Tr}[\Pi_{i}]+e^{-\beta\lambda_{1}}\operatorname{Tr}[\Pi_{1}]}$
$\qquad+1-\frac{e^{-\beta\lambda_{1}}\operatorname{Tr}[\Pi_{1}]}{\sum_{i=1}^{M}e^{-\beta\lambda_{i}}\operatorname{Tr}[\Pi_{i}]}$ (156)
$=2\frac{\sum_{i=2}^{M}e^{-\beta\lambda_{i}}\operatorname{Tr}[\Pi_{i}]}{\sum_{i=2}^{M}e^{-\beta\lambda_{i}}\operatorname{Tr}[\Pi_{i}]+e^{-\beta\lambda_{1}}\operatorname{Tr}[\Pi_{1}]}$ (157)
$=2\frac{\sum_{i=2}^{M}e^{-\beta\left(\lambda_{i}-\lambda_{1}\right)}\operatorname{Tr}[\Pi_{i}]}{\sum_{i=2}^{M}e^{-\beta\left(\lambda_{i}-\lambda_{1}\right)}\operatorname{Tr}[\Pi_{i}]+\operatorname{Tr}[\Pi_{1}]}$ (158)
$=2\frac{1}{1+\frac{\operatorname{Tr}[\Pi_{1}]}{\sum_{i=2}^{M}e^{-\beta\left(\lambda_{i}-\lambda_{1}\right)}\operatorname{Tr}[\Pi_{i}]}}.$ (159)
Now recall that $\Delta=\lambda_{2}-\lambda_{1}$. Then $e^{-\beta\left(\lambda_{i}-\lambda_{1}\right)}\leq e^{\beta\Delta}$ for all $i\in\left\{2,\ldots,M\right\}$, which implies that
$\frac{1}{1+\frac{\operatorname{Tr}[\Pi_{1}]}{\sum_{i=2}^{M}e^{-\beta\left(\lambda_{i}-\lambda_{1}\right)}\operatorname{Tr}[\Pi_{i}]}}$
$\leq\frac{1}{1+\frac{\operatorname{Tr}[\Pi_{1}]}{\sum_{i=2}^{M}e^{-\beta\Delta}\operatorname{Tr}[\Pi_{i}]}}$ (160)
$=\frac{1}{1+e^{\beta\Delta}\frac{\operatorname{Tr}[\Pi_{1}]}{\sum_{i=2}^{M}\operatorname{Tr}[\Pi_{i}]}}$ (161)
$=\frac{1}{1+e^{\beta\Delta}\frac{\operatorname{Tr}[\Pi_{1}]}{\operatorname{Tr}\left[\sum_{i=2}^{M}\Pi_{i}\right]}}$ (162)
$=\frac{1}{1+e^{\beta\Delta}\frac{\operatorname{Tr}[\Pi_{1}]}{\operatorname{Tr}\left[I-\Pi_{1}\right]}}$ (163)
$=\frac{1}{1+e^{\beta\Delta}\frac{\operatorname{Tr}[\Pi_{1}]}{d-\operatorname{Tr}\left[\Pi_{1}\right]}},$ (164)
thus concluding the proof. ∎
Corollary 10.
For the normalized trace distance to be $\leq\varepsilon\in[0,1]$ in (147), i.e., to have
$\frac{1}{2}\left\|\frac{e^{-\beta H}}{\operatorname{Tr}[e^{-\beta H}]}-\frac{\Pi_{1}}{\operatorname{Tr}[\Pi_{1}]}\right\|_{1}\leq\varepsilon,$ (165)
we require the inverse temperature $\beta$ to be set as follows:
$\beta=\frac{1}{\Delta}\ln\!\left[\left(\frac{1-\varepsilon}{\varepsilon}\right)\left(\frac{d-d_{G}}{d_{G}}\right)\right],$ (166)
Proof.
Consider that we need to solve for $\varepsilon$ in (167) and find that
$\frac{1}{1+e^{\beta\Delta}\frac{\operatorname{Tr}[\Pi_{1}]}{d-\operatorname{Tr}\left[\Pi_{1}\right]}}$ $=\varepsilon$ (167)
$\iff\quad 1+e^{\beta\Delta}\frac{\operatorname{Tr}[\Pi_{1}]}{d-\operatorname{Tr}\left[\Pi_{1}\right]}$ $=\frac{1}{\varepsilon}$ (168)
$\iff\quad e^{\beta\Delta}\frac{\operatorname{Tr}[\Pi_{1}]}{d-\operatorname{Tr}\left[\Pi_{1}\right]}$ $=\frac{1}{\varepsilon}-1$ (169)
$\iff\quad\left(\frac{1-\varepsilon}{\varepsilon}\right)\left(\frac{d-\operatorname{Tr}\left[\Pi_{1}\right]}{\operatorname{Tr}[\Pi_{1}]}\right)$ $=e^{\beta\Delta}$ (170)
$\iff\quad\frac{1}{\Delta}\ln\!\left[\left(\frac{1-\varepsilon}{\varepsilon}\right)\left(\frac{d-\operatorname{Tr}\left[\Pi_{1}\right]}{\operatorname{Tr}[\Pi_{1}]}\right)\right]=\beta,$ (171)
concluding the proof. ∎
Proposition 11.
For a $d$-dimensional Hamiltonian $H$ as in (145) and $\beta\geq 0$, the following hold:
$F\!\left(\frac{e^{-\beta H}}{\operatorname{Tr}[e^{-\beta H}]},\frac{\Pi_{1}}{\operatorname{Tr}[\Pi_{1}]}\right)$
$=\frac{1}{1+\frac{\sum_{i=2}^{M}e^{-\beta(\lambda_{i}-\lambda_{1})}\operatorname{Tr}[\Pi_{i}]}{d_{G}}}$ (172)
$\geq\frac{1}{1+e^{-\beta\Delta}\left(\frac{d-d_{G}}{d_{G}}\right)},$ (173)
where we have used the same notation as in Proposition 9.
Proof.
Consider that
$F\!\left(\frac{e^{-\beta H}}{\operatorname{Tr}[e^{-\beta H}]},\frac{\Pi_{1}}{\operatorname{Tr}[\Pi_{1}]}\right)$
$=\left(\operatorname{Tr}\!\left[\sqrt{\sqrt{\frac{\Pi_{1}}{\operatorname{Tr}[\Pi_{1}]}}\frac{e^{-\beta H}}{\operatorname{Tr}[e^{-\beta H}]}\sqrt{\frac{\Pi_{1}}{\operatorname{Tr}[\Pi_{1}]}}}\right]\right)^{2}$ (174)
$=\frac{\left(\operatorname{Tr}\!\left[\sqrt{\Pi_{1}e^{-\beta H}\Pi_{1}}\right]\right)^{2}}{\operatorname{Tr}[\Pi_{1}]\operatorname{Tr}[e^{-\beta H}]}$ (175)
$=\frac{\left(\operatorname{Tr}\!\left[\sqrt{\Pi_{1}\left(\sum_{i=1}^{M}e^{-\beta\lambda_{i}}\Pi_{i}\right)\Pi_{1}}\right]\right)^{2}}{\operatorname{Tr}[\Pi_{1}]\sum_{i=1}^{M}e^{-\beta\lambda_{i}}\operatorname{Tr}[\Pi_{i}]}$ (176)
$=\frac{\left(\operatorname{Tr}\!\left[\sqrt{e^{-\beta\lambda_{1}}\Pi_{1}}\right]\right)^{2}}{\operatorname{Tr}[\Pi_{1}]\sum_{i=1}^{M}e^{-\beta\lambda_{i}}\operatorname{Tr}[\Pi_{i}]}$ (177)
$=\frac{e^{-\beta\lambda_{1}}\operatorname{Tr}[\Pi_{1}]}{\sum_{i=1}^{M}e^{-\beta\lambda_{i}}\operatorname{Tr}[\Pi_{i}]}$ (178)
$=\frac{e^{-\beta\lambda_{1}}\operatorname{Tr}[\Pi_{1}]}{e^{-\beta\lambda_{1}}\operatorname{Tr}[\Pi_{1}]+\sum_{i=2}^{M}e^{-\beta\lambda_{i}}\operatorname{Tr}[\Pi_{i}]}$ (179)
$=\frac{1}{1+\frac{1}{d_{G}}\sum_{i=2}^{M}e^{-\beta(\lambda_{i}-\lambda_{1})}\operatorname{Tr}[\Pi_{i}]}$ (180)
$\geq\frac{\operatorname{Tr}[\Pi_{1}]}{1+\frac{1}{d_{G}}\sum_{i=2}^{M}e^{-\beta\Delta}\operatorname{Tr}[\Pi_{i}]}$ (181)
$=\frac{1}{1+\frac{e^{-\beta\Delta}}{d_{G}}\sum_{i=2}^{M}\operatorname{Tr}[\Pi_{i}]}$ (182)
$=\frac{1}{1+e^{-\beta\Delta}\left(\frac{d-d_{G}}{d_{G}}\right)}.$ (183)
This completes the proof. ∎
Corollary 12.
For the fidelity to be $\geq 1-\varepsilon$ in (173) for $\varepsilon\in[0,1]$, i.e., to have
$F\!\left(\frac{e^{-\beta H}}{\operatorname{Tr}[e^{-\beta H}]},\frac{\Pi_{1}}{\operatorname{Tr}[\Pi_{1}]}\right)\geq 1-\varepsilon,$ (184)
we require the inverse temperature $\beta$ to be set as follows:
$\beta=\frac{1}{\Delta}\ln\!\left[\left(\frac{1-\varepsilon}{\varepsilon}\right)\left(\frac{d-d_{G}}{d_{G}}\right)\right].$ (185)
Proof.
Considering (173), we should solve for $\beta$ therein and find that
$\frac{1}{1+e^{-\beta\Delta}\left(\frac{d-d_{G}}{d_{G}}\right)}$ $=1-\varepsilon$ (186)
$\iff\quad 1+e^{-\beta\Delta}\left(\frac{d-d_{G}}{d_{G}}\right)$ $=\frac{1}{1-\varepsilon}$ (187)
$\iff\quad e^{-\beta\Delta}\left(\frac{d-d_{G}}{d_{G}}\right)$ $=\frac{1}{1-\varepsilon}-1$ (188)
$\iff\quad e^{-\beta\Delta}$ $=\left(\frac{\varepsilon}{1-\varepsilon}\right)\left(\frac{d_{G}}{d-d_{G}}\right)$ (189)
$\iff\quad\beta=\frac{1}{\Delta}\ln\!\left[\left(\frac{1-\varepsilon}{\varepsilon}\right)\left(\frac{d-d_{G}}{d_{G}}\right)\right],$ (190)
thus concluding the proof. ∎
Proposition 13.
For a $d$-dimensional Hamiltonian $H$ as in (145) and $\beta\geq 0$, the following hold:
$D\!\left(\frac{\Pi_{1}}{\operatorname{Tr}[\Pi_{1}]}\middle\|\frac{e^{-\beta H}}{\operatorname{Tr}[e^{-\beta H}]}\right)$
$=\ln\!\left(1+\frac{\sum_{i=2}^{M}e^{-\beta\left(\lambda_{i}-\lambda_{1}\right)}\operatorname{Tr}[\Pi_{i}]}{d_{G}}\right)$ (191)
$\leq\ln\!\left(1+e^{-\beta\Delta}\left(\frac{d-d_{G}}{d_{G}}\right)\right),$ (192)
where we have used the same notation as in Proposition 9. Furthermore, the following equalities hold for all $\alpha\in(0,1)\cup(1,\infty)$:
$D\!\left(\frac{\Pi_{1}}{\operatorname{Tr}[\Pi_{1}]}\middle\|\frac{e^{-\beta H}}{\operatorname{Tr}[e^{-\beta H}]}\right)$
$=D_{\alpha}\!\left(\frac{\Pi_{1}}{\operatorname{Tr}[\Pi_{1}]}\middle\|\frac{e^{-\beta H}}{\operatorname{Tr}[e^{-\beta H}]}\right)$ (193)
$=\widetilde{D}_{\alpha}\!\left(\frac{\Pi_{1}}{\operatorname{Tr}[\Pi_{1}]}\middle\|\frac{e^{-\beta H}}{\operatorname{Tr}[e^{-\beta H}]}\right)$ (194)
$=\widehat{D}_{\alpha}\!\left(\frac{\Pi_{1}}{\operatorname{Tr}[\Pi_{1}]}\middle\|\frac{e^{-\beta H}}{\operatorname{Tr}[e^{-\beta H}]}\right).$ (195)
Proof.
Consider that
$D\!\left(\frac{\Pi_{1}}{\operatorname{Tr}[\Pi_{1}]}\middle\|\frac{e^{-\beta H}}{\operatorname{Tr}[e^{-\beta H}]}\right)$
$=\operatorname{Tr}\!\left[\frac{\Pi_{1}}{\operatorname{Tr}[\Pi_{1}]}\ln\!\left(\frac{\Pi_{1}}{\operatorname{Tr}[\Pi_{1}]}\right)\right]$
$\qquad-\operatorname{Tr}\!\left[\frac{\Pi_{1}}{\operatorname{Tr}[\Pi_{1}]}\ln\!\left(\frac{e^{-\beta H}}{\operatorname{Tr}[e^{-\beta H}]}\right)\right]$ (196)
$=-\ln d_{G}-\operatorname{Tr}\!\left[\frac{\Pi_{1}}{\operatorname{Tr}[\Pi_{1}]}\ln\!\left(e^{-\beta H}\right)\right]+\ln\operatorname{Tr}[e^{-\beta H}]$ (197)
$=-\ln d_{G}+\frac{\beta}{\operatorname{Tr}[\Pi_{1}]}\operatorname{Tr}\!\left[\Pi_{1}H\right]+\ln\operatorname{Tr}[e^{-\beta H}]$ (198)
$=-\ln d_{G}+\frac{\lambda_{1}\beta}{\operatorname{Tr}[\Pi_{1}]}\operatorname{Tr}\!\left[\Pi_{1}\right]+\ln\operatorname{Tr}[e^{-\beta H}]$ (199)
$=\lambda_{1}\beta-\ln d_{G}+\ln\operatorname{Tr}[e^{-\beta H}]$ (200)
$=\lambda_{1}\beta-\ln d_{G}+\ln\sum_{i=1}^{M}e^{-\beta\lambda_{i}}\operatorname{Tr}[\Pi_{i}]$ (201)
$=\ln\!\left(\frac{1}{d_{G}}\sum_{i=1}^{M}e^{-\beta\left(\lambda_{i}-\lambda_{1}\right)}\operatorname{Tr}[\Pi_{i}]\right)$ (202)
$=\ln\!\left(\frac{\operatorname{Tr}[\Pi_{1}]+\sum_{i=2}^{M}e^{-\beta\left(\lambda_{i}-\lambda_{1}\right)}\operatorname{Tr}[\Pi_{i}]}{d_{G}}\right)$ (203)
$=\ln\!\left(1+\frac{1}{d_{G}}\sum_{i=2}^{M}e^{-\beta\left(\lambda_{i}-\lambda_{1}\right)}\operatorname{Tr}[\Pi_{i}]\right)$ (204)
$\leq\ln\!\left(1+\frac{1}{d_{G}}\sum_{i=2}^{M}e^{-\beta\Delta}\operatorname{Tr}[\Pi_{i}]\right)$ (205)
$=\ln\!\left(1+e^{-\beta\Delta}\left(\frac{d-d_{G}}{d_{G}}\right)\right).$ (206)
This establishes (191) and (192).
We now prove (193)–(195). Since the states of interest commute, the equalities in (144) follow, and it suffices to evaluate the Petz–Rényi relative entropy. To this end, consider that
$D_{\alpha}\!\left(\frac{\Pi_{1}}{\operatorname{Tr}[\Pi_{1}]}\middle\|\frac{e^{-\beta H}}{\operatorname{Tr}[e^{-\beta H}]}\right)$
$=\frac{1}{\alpha-1}\ln\operatorname{Tr}\!\left[\left(\frac{\Pi_{1}}{\operatorname{Tr}[\Pi_{1}]}\right)^{\alpha}\left(\frac{e^{-\beta H}}{\operatorname{Tr}[e^{-\beta H}]}\right)^{1-\alpha}\right]$ (207)
$=\frac{1}{\alpha-1}\ln\frac{1}{d_{G}^{\alpha}\left(\operatorname{Tr}[e^{-\beta H}]\right)^{1-\alpha}}$
$\qquad+\frac{1}{\alpha-1}\ln\operatorname{Tr}\!\left[\Pi_{1}^{\alpha}e^{-\beta\left(1-\alpha\right)H}\right]$ (208)
$=-\frac{\alpha}{\alpha-1}\ln d_{G}+\ln\operatorname{Tr}[e^{-\beta H}]$
$\qquad+\frac{1}{\alpha-1}\ln\operatorname{Tr}\!\left[e^{-\beta\left(1-\alpha\right)\lambda_{1}}\Pi_{1}\right]$ (209)
$=-\frac{\alpha}{\alpha-1}\ln d_{G}+\ln\operatorname{Tr}[e^{-\beta H}]$
$\qquad+\beta\lambda_{1}+\frac{1}{\alpha-1}\ln d_{G}$ (210)
$=-\ln d_{G}+\ln\operatorname{Tr}[e^{-\beta H}]+\beta\lambda_{1}.$ (211)
This last expression coincides with that in (200), thus concluding the proof. ∎
Corollary 14.
Fix $\varepsilon>0$. For the relative entropy to be $\leq\varepsilon$ in (192), i.e., to have
$D\!\left(\frac{\Pi_{1}}{\operatorname{Tr}[\Pi_{1}]}\middle\|\frac{e^{-\beta H}}{\operatorname{Tr}[e^{-\beta H}]}\right)\leq\varepsilon,$ (212)
we require the inverse temperature $\beta$ to be set as follows:
$\beta=\frac{1}{\Delta}\ln\!\left[\left(\frac{1}{e^{\varepsilon}-1}\right)\left(\frac{d-d_{G}}{d_{G}}\right)\right].$ (213)
Due to (193)–(195), the same statement applies to the Petz, sandwiched, and geometric Rényi relative entropies for all $\alpha\in(0,1)\cup(1,\infty)$.
Proof.
We have to solve for the inverse temperature $\beta$ in the following equation:
$\ln\!\left(1+e^{-\beta\Delta}\left(\frac{d-d_{G}}{d_{G}}\right)\right)$ $=\varepsilon$ (214)
$\iff\qquad e^{\varepsilon}-1$ $=e^{-\beta\Delta}\left(\frac{d-d_{G}}{d_{G}}\right)$ (215)
$\iff\qquad\frac{d_{G}}{d-d_{G}}\left(e^{\varepsilon}-1\right)$ $=e^{-\beta\Delta}$ (216)
$\iff\qquad\beta=\frac{1}{\Delta}\ln\!\left[\left(\frac{1}{e^{\varepsilon}-1}\right)\left(\frac{d-d_{G}}{d_{G}}\right)\right],$ (217)
thus concluding the proof. ∎
The following corollary establishes equalities relating several distinguishability measures considered in this appendix, whenever the two states being compared are the temperature-$T$ thermal state of a Hamiltonian and the maximally mixed state on its ground space.
Corollary 15.
For a $d$-dimensional Hamiltonian $H$ as in (145) and $\beta\geq 0$, the following equalities hold:
$\frac{1}{2}\left\|\frac{e^{-\beta H}}{\operatorname{Tr}[e^{-\beta H}]}-\frac{\Pi_{1}}{\operatorname{Tr}[\Pi_{1}]}\right\|_{1}=\\ 1-F\!\left(\frac{e^{-\beta H}}{\operatorname{Tr}[e^{-\beta H}]},\frac{\Pi_{1}}{\operatorname{Tr}[\Pi_{1}]}\right),$ (218)
$D\!\left(\frac{\Pi_{1}}{\operatorname{Tr}[\Pi_{1}]}\middle\|\frac{e^{-\beta H}}{\operatorname{Tr}[e^{-\beta H}]}\right)=\\ -\ln\!\left(F\!\left(\frac{e^{-\beta H}}{\operatorname{Tr}[e^{-\beta H}]},\frac{\Pi_{1}}{\operatorname{Tr}[\Pi_{1}]}\right)\right)$ (219)
Proof.
Set
$a$ $\equiv\frac{d_{G}}{\sum_{i=2}^{M}e^{-\beta\left(\lambda_{i}-\lambda_{1}\right)}\operatorname{Tr}[\Pi_{i}]},$ (220)
$T$ $\equiv\frac{1}{2}\left\|\frac{e^{-\beta H}}{\operatorname{Tr}[e^{-\beta H}]}-\frac{\Pi_{1}}{\operatorname{Tr}[\Pi_{1}]}\right\|_{1},$ (221)
$F$ $\equiv F\!\left(\frac{e^{-\beta H}}{\operatorname{Tr}[e^{-\beta H}]},\frac{\Pi_{1}}{\operatorname{Tr}[\Pi_{1}]}\right),$ (222)
$D$ $\equiv D\!\left(\frac{\Pi_{1}}{\operatorname{Tr}[\Pi_{1}]}\middle\|\frac{e^{-\beta H}}{\operatorname{Tr}[e^{-\beta H}]}\right).$ (223)
By Propositions 9, 11, and 13, we have that
$T$ $=\frac{1}{1+a},$ (224)
$F$ $=\frac{1}{1+\frac{1}{a}},$ (225)
$D$ $=\ln\!\left(1+\frac{1}{a}\right).$ (226)
Clearly, $D=-\ln F$. Additionally,
$1-F$ $=1-\frac{1}{1+\frac{1}{a}}=\frac{1+\frac{1}{a}}{1+\frac{1}{a}}-\frac{1}{1+\frac{1}{a}}$ (227)
$=\frac{\frac{1}{a}}{1+\frac{1}{a}}=\frac{1}{1+a}$ (228)
$=T,$ (229)
thus concluding the proof. ∎
Appendix B Simplifying the implementation of the channel $\Phi_{\mu}$ for extensive, conserved charges
In this appendix, we provide a proof of (39). Starting from (37), Eq. (39) follows because
$\Phi_{\mu}(Q_{i})$
$=\int_{-\infty}^{\infty}dt\ p(t)\,e^{i\mu\cdot Qt/T}Q_{i}e^{-i\mu\cdot Qt/T}$ (230)
$=\int_{-\infty}^{\infty}dt\ p(t)\,e^{i\mu\cdot Qt/T}\left(\sum_{j=1}^{n}C_{i}^{(j)}\right)e^{-i\mu\cdot Qt/T}$ (231)
$=\sum_{j=1}^{n}\int_{-\infty}^{\infty}dt\ p(t)\,e^{i\mu\cdot Qt/T}C_{i}^{(j)}e^{-i\mu\cdot Qt/T}$ (232)
$=\sum_{j=1}^{n}\int_{-\infty}^{\infty}dt\ p(t)\,e^{\frac{it}{T}\sum_{i^{\prime}\in[c]}\mu_{i^{\prime}}Q_{i^{\prime}}}C_{i}^{(j)}e^{-\frac{it}{T}\sum_{i^{\prime\prime}\in[c]}\mu_{i^{\prime\prime}}Q_{i^{\prime\prime}}}$ (233)
$=\sum_{j=1}^{n}\int_{-\infty}^{\infty}dt\ p(t)\,e^{\frac{it}{T}\sum_{i^{\prime}\in[c]}\mu_{i^{\prime}}\sum_{j^{\prime}=1}^{n}C_{i^{\prime}}^{(j^{\prime})}}C_{i}^{(j)}\times$
$\qquad\qquad\qquad\qquad e^{-\frac{it}{T}\sum_{i^{\prime\prime}\in[c]}\mu_{i^{\prime\prime}}\sum_{j^{\prime\prime}=1}^{n}C_{i^{\prime\prime}}^{(j^{\prime\prime})}}$ (234)
$=\sum_{j=1}^{n}\int_{-\infty}^{\infty}dt\ p(t)\,\prod_{j^{\prime}=1}^{n}e^{\frac{it}{T}\sum_{i^{\prime}\in[c]}\mu_{i^{\prime}}C_{i^{\prime}}^{(j^{\prime})}}C_{i}^{(j)}\times$
$\qquad\qquad\qquad\qquad\prod_{j^{\prime\prime}=1}^{n}e^{-\frac{it}{T}\sum_{i^{\prime\prime}\in[c]}\mu_{i^{\prime\prime}}C_{i^{\prime\prime}}^{(j^{\prime\prime})}}$ (235)
$=\sum_{j=1}^{n}\int_{-\infty}^{\infty}dt\ p(t)\,e^{\frac{it}{T}\sum_{i^{\prime}\in[c]}\mu_{i^{\prime}}C_{i^{\prime}}^{(j)}}C_{i}^{(j)}\times$
$\qquad\qquad\qquad\qquad e^{-\frac{it}{T}\sum_{i^{\prime\prime}\in[c]}\mu_{i^{\prime\prime}}C_{i^{\prime\prime}}^{(j)}}$ (236)
$=\sum_{j=1}^{n}\int_{-\infty}^{\infty}dt\ p(t)\,e^{\frac{it}{T}\mu\cdot C^{(j)}}C_{i}^{(j)}e^{-\frac{it}{T}\mu\cdot C^{(j)}}.$ (237)
Eqs. (231) and (234) follow by substituting (5). Eqs. (235) and (236) follow by observing that $\left[C_{i^{\prime}}^{(j^{\prime})},C_{i^{\prime\prime}}^{(j^{\prime\prime})}\right]=0$ for all $j^{\prime}\neq j^{\prime\prime}$ and applying (38).
Appendix C Derivation of Equation (64)
In this appendix, we include a derivation of Eq. (64) for completeness.
First, recall that a Clifford unitary is one that preserves the $n$-fold Pauli group $\mathcal{P}_{n}$ when acting on its elements by conjugation [104] (see also [39, Section 3.5]); i.e., the set $C_{n}$ of all Clifford unitaries acting on $n$ qubits is defined as follows:
$C_{n}\coloneqq\left\{U\in\mathcal{U}(2^{n}):UPU^{{\dagger}}\in\mathcal{P}_{n}\,\forall P\in\mathcal{P}_{n}\right\},$ (238)
where $\mathcal{U}(2^{n})$ denotes the set of all $n$-qubit unitaries.
Consider a stabilizer code that encodes $k\in\mathbb{N}$ logical qubits into $n\in\mathbb{N}$ physical qubits, where $k\leq n$. Such a code is encoded by a Clifford encoding unitary $U$. The unencoded state before the encoded unitary acts is as follows:
$\rho\otimes|0\rangle\!\langle 0|^{\otimes n-k},$ (239)
where $\rho$ is defined as in (63), and the state after the encoding is as follows:
$U\left(\rho\otimes|0\rangle\!\langle 0|^{\otimes n-k}\right)U^{{\dagger}}.$ (240)
Recalling that $|0\rangle\!\langle 0|=\frac{I+Z}{2}$, we can write the state before encoding takes place as follows:
$\rho\otimes\left(\frac{I+Z}{2}\right)^{\otimes n-k}\\ =\left(I^{\otimes k}\otimes\left(\frac{I+Z}{2}\right)^{\otimes n-k}\right)\left(\rho\otimes I^{\otimes n-k}\right)\times\\ \left(I^{\otimes k}\otimes\left(\frac{I+Z}{2}\right)^{\otimes n-k}\right),$ (241)
so that the state after encoding takes place is
$U\left[\rho\otimes\left(\frac{I+Z}{2}\right)^{\otimes n-k}\right]U^{{\dagger}}\\ =\left[U\left(I^{\otimes k}\otimes\left(\frac{I+Z}{2}\right)^{\otimes n-k}\right)U^{{\dagger}}\right]\times\\ \left[U\left(\rho\otimes I^{\otimes n-k}\right)U^{{\dagger}}\right]\times\\ \left[U\left(I^{\otimes k}\otimes\left(\frac{I+Z}{2}\right)^{\otimes n-k}\right)U^{{\dagger}}\right].$ (242)
Now recall that the action of the Clifford unitary is to transform the nkn-k operators $Z_{k+1},\ldots,Z_{n}$ to nkn-k stabilizer generators $\overline{Z}_{k+1},\ldots,\overline{Z}_{n}$ as follows:
$UZ_{i}U^{{\dagger}}=\overline{Z}_{i}\qquad\forall i\in\left\{k+1,\ldots,n\right\},$ (243)
and it transforms the $3k$ logical operators $X_{1}$, $Y_{1}$, $Z_{1}$, …, $X_{k}$, $Y_{k}$, $Z_{k}$ for the unencoded state to the $3k$ logical operators $\overline{X}_{1}$, $\overline{Y}_{1}$, $\overline{Z}_{1}$, …, $\overline{X}_{k}$, $\overline{Y}_{k}$, $\overline{Z}_{k}$ for the encoded state, as follows:
$UX_{i}U^{{\dagger}}$ $=\overline{X}_{i}\qquad\forall i\in\left[k\right],$ (244)
$UY_{i}U^{{\dagger}}$ $=\overline{Y}_{i}\qquad\forall i\in\left[k\right],$ (245)
$UZ_{i}U^{{\dagger}}$ $=\overline{Z}_{i}\qquad\forall i\in\left[k\right].$ (246)
As such, it follows that
$U\left(I^{\otimes k}\otimes\left(\frac{I+Z}{2}\right)^{\otimes n-k}\right)U^{{\dagger}}$
$=U\left(\prod_{i=k+1}^{n}\frac{I+Z_{i}}{2}\right)U^{{\dagger}}$ (247)
$=\prod_{i=k+1}^{n}U\left(\frac{I+Z_{i}}{2}\right)U^{{\dagger}}$ (248)
$=\prod_{i=k+1}^{n}\left(\frac{I+\overline{Z}_{i}}{2}\right)$ (249)
$=\Pi_{\mathcal{C}},$ (250)
and
$U\left(\rho\otimes I^{\otimes n-k}\right)U^{{\dagger}}=\\ \frac{1}{2^{k}}\sum_{i_{1},\ldots,i_{k}\in\{0,1,2,3\}}r_{i_{1},\ldots,i_{k}}\overline{\sigma}_{i_{1},1}\cdots\overline{\sigma}_{i_{k},k},$ (251)
where we have applied the definitions in (63) and (64). Thus, combining (242) and (247)–(251), we conclude (64).
Appendix D Derivation of Equation (77)
Let us redo the development in (71)–(75) in a slightly different way, in order to arrive at the complementary slackness condition in (77). Let us set $Q_{0}=I$ and $q_{0}=1$, in order to encode the constraint that $\rho$ is a density matrix, i.e., $\operatorname{Tr}[\rho]=1$. Consider that
$\min_{\rho\in\mathcal{D}_{d^{n}}}\left\{\operatorname{Tr}[H\rho]:\operatorname{Tr}[Q_{i}\rho]=q_{i}\ \forall i\in\left[c\right]\right\}$
$=\min_{\rho\geq 0}\left\{\operatorname{Tr}[H\rho]:\operatorname{Tr}[Q_{i}\rho]=q_{i}\ \forall i\in\left\{0,1,\ldots,c\right\}\right\}$ (252)
$=\min_{\rho\geq 0}\left\{\operatorname{Tr}[H\rho]+\sup_{\begin{subarray}{c}\mu\in\mathbb{R}^{c},\\ \mu_{0}\in\mathbb{R}\end{subarray}}\sum_{i\in\left\{0,1,\ldots,c\right\}}\mu_{i}\left(q_{i}-\operatorname{Tr}[Q_{i}\rho]\right)\right\}$ (253)
$=\min_{\rho\geq 0}\sup_{\begin{subarray}{c}\mu\in\mathbb{R}^{c},\\ \mu_{0}\in\mathbb{R}\end{subarray}}\left\{\operatorname{Tr}[H\rho]+\sum_{i\in\left\{0,1,\ldots,c\right\}}\mu_{i}\left(q_{i}-\operatorname{Tr}[Q_{i}\rho]\right)\right\}$ (254)
$=\min_{\rho\geq 0}\sup_{\begin{subarray}{c}\mu\in\mathbb{R}^{c},\\ \mu_{0}\in\mathbb{R}\end{subarray}}\left\{\mu\cdot q+\mu_{0}+\operatorname{Tr}[\left(H-\mu\cdot Q-\mu_{0}I\right)\rho]\right\}$ (255)
$=\sup_{\begin{subarray}{c}\mu\in\mathbb{R}^{c},\\ \mu_{0}\in\mathbb{R}\end{subarray}}\min_{\rho\geq 0}\left\{\mu\cdot q+\mu_{0}+\operatorname{Tr}[\left(H-\mu\cdot Q-\mu_{0}I\right)\rho]\right\}$ (256)
$=\sup_{\begin{subarray}{c}\mu\in\mathbb{R}^{c},\\ \mu_{0}\in\mathbb{R}\end{subarray}}\left\{\mu\cdot q+\mu_{0}:H-\mu\cdot Q\geq\mu_{0}I\right\}.$ (257)
The equalities follow for reasons similar to those given for (71)–(75), with the exception that (256) follows from strong duality of semi-definite programming, under the assumption that there exists at least one state $\rho$ satisfying the constraints in (69). Observe that an optimal $\mu_{0}$ in (257) is equal to the minimum eigenvalue of $H-\mu\cdot Q$. Now suppose that $\rho^{*}$ is optimal for (252) and $\mu^{*}$ is optimal for (257). Then it follows that
$\operatorname{Tr}[H\rho^{*}]$ $=\mu^{*}\cdot q+\mu_{0}$ (258)
$=\mu^{*}\cdot q+\lambda_{\min}(H-\mu^{*}\cdot Q).$ (259)
Consider that $\operatorname{Tr}[Q_{i}\rho^{*}]=q_{i}$ for all $i\in\left[c\right]$. Then
$\mu^{*}\cdot q=\operatorname{Tr}[\left(\mu^{*}\cdot Q\right)\rho^{*}].$ (260)
Plugging back in above, we conclude the desired equality:
$\operatorname{Tr}[H\rho^{*}]$ $=\operatorname{Tr}[\left(\mu^{*}\cdot Q\right)\rho]+\lambda_{\min}(H-\mu^{*}\cdot Q)$ (261)
$\iff\qquad$ $\operatorname{Tr}[\left(H-\mu^{*}\cdot Q\right)\rho^{*}]=\lambda_{\min}(H-\mu^{*}\cdot Q),$ (262)
so that $\rho^{*}$ is in the ground space of $H-\mu^{*}\cdot Q$.
Appendix E Proof of Theorem 7
Proof.
Consider that, in order for the equality in (115) to hold, the eigenvectors of the matrix on the left-hand side and those for the matrix on the right-hand side should match. The eigenvectors for the matrix on the left-hand side are the same as those for the following matrix:
$r_{x}X+r_{y}Y+r_{z}Z,$ (263)
and the eigenvectors for the matrix on the right-hand side are the same as those for the following matrix:
$\mu_{x}X+\mu_{y}Y+\mu_{z}Z.$ (264)
Thus, for the eigenvectors to match, we simply set $\mu=r$. This is also intuitive, because the vector $r$ identifies a direction in the Bloch sphere, and so does $\mu$, and these directions should be aligned. So now we are trying to solve the following equation for $\beta$, which is the only parameter left to solve for:
$\frac{1}{2}\left(I+r_{x}X+r_{y}Y+r_{z}Z\right)=\\ \frac{1}{Z_{\beta}(r)}\exp\!\left(-\beta\left(r_{x}X+r_{y}Y+r_{z}Z\right)\right).$ (265)
It is well known that the eigenvalues of the matrix in (263) are given by
$\pm\left\|r\right\|.$ (266)
To see this, recall that the eigenvalues of a $2\times 2$ matrix $A$ are given by the following formula:
$\lambda_{\pm}=\frac{1}{2}\left(\operatorname{Tr}[A]\pm\sqrt{\left(\operatorname{Tr}[A]\right)^{2}-4\det(A)}\right).$ (267)
Since the trace of the matrix in (263) is equal to zero, this implies that the eigenvalues are given by the following formula:
$\pm\sqrt{-\det(r_{x}X+r_{y}Y+r_{z}Z)},$ (268)
which one finds upon direct calculation to be equal to (266). This implies that the eigenvalues of the matrix in (110) are given by
$\frac{1\pm\left\|r\right\|}{2},$ (269)
and the eigenvalues of $\exp\!\left(-\beta\left(r_{x}X+r_{y}Y+r_{z}Z\right)\right)$ are given by
$e^{\mp\beta\left\|r\right\|}.$ (270)
Then
$Z_{\beta}(r)=e^{-\beta\left\|r\right\|}+e^{\beta\left\|r\right\|},$ (271)
and we conclude that the eigenvalues of the matrix on the right-hand side of (265) are given by
$\frac{e^{\mp\beta\left\|r\right\|}}{e^{-\beta\left\|r\right\|}+e^{\beta\left\|r\right\|}}.$ (272)
Now setting the first eigenvalue in (269) to the first eigenvalue in (272), we conclude that
$\frac{1+\left\|r\right\|}{2}=\frac{e^{-\beta\left\|r\right\|}}{e^{-\beta\left\|r\right\|}+e^{\beta\left\|r\right\|}}.$ (273)
Since $\beta$ is a free parameter, let us set
$\gamma=\beta\left\|r\right\|$ (274)
and rewrite (273) as follows:
$\frac{1+\left\|r\right\|}{2}=\frac{e^{-\gamma}}{e^{-\gamma}+e^{\gamma}}.$ (275)
Now let us solve this equation for $\gamma$:
$\frac{1+\left\|r\right\|}{2}$ $=\frac{e^{-\gamma}}{e^{-\gamma}+e^{\gamma}}$ (276)
$\iff\qquad$ $1+\left\|r\right\|=\frac{2e^{-\gamma}}{e^{-\gamma}+e^{\gamma}}$ (277)
$\iff\qquad$ $\left\|r\right\|=\frac{2e^{-\gamma}}{e^{-\gamma}+e^{\gamma}}-\frac{e^{-\gamma}+e^{\gamma}}{e^{-\gamma}+e^{\gamma}}$ (278)
$\iff\qquad$ $\left\|r\right\|=\frac{e^{-\gamma}-e^{\gamma}}{e^{-\gamma}+e^{\gamma}}$ (279)
$\iff\qquad$ $\left\|r\right\|=-\tanh(\gamma).$ (280)
This last line implies that
$\gamma=\operatorname{arctanh}\!\left(-\left\|r\right\|\right).$ (281)
Finally, substituting into (274), we find that
$\beta=\frac{\operatorname{arctanh}\!\left(-\left\|r\right\|\right)}{\left\|r\right\|}.$ (282)
This concludes our first proof.
We also provide a second, alternative proof based on Taylor expansions. Consider that the following equality should hold:
$\frac{1}{2}\left(I+\overrightarrow{r}\cdot\overrightarrow{\sigma}\right)=\frac{\exp\!\left(-\beta\overrightarrow{\mu}\cdot\overrightarrow{\sigma}\right)}{\operatorname{Tr}\!\left[\exp\!\left(-\beta\overrightarrow{\mu}\cdot\overrightarrow{\sigma}\right)\right]},$ (283)
where
$\overrightarrow{r}$ $\coloneqq\left(r_{x},r_{y},r_{z}\right),$ (284)
$\overrightarrow{\sigma}$ $\coloneqq\left(\sigma_{X},\sigma_{Y},\sigma_{Z}\right),$ (285)
$\overrightarrow{\mu}$ $\coloneqq\left(\mu_{x},\mu_{y},\mu_{z}\right).$ (286)
Without loss of generality, let us suppose that $\overrightarrow{\mu}$ is normalized, so that $\left\|\overrightarrow{\mu}\right\|=1$. Observe that $\left(\overrightarrow{\mu}\cdot\overrightarrow{\sigma}\right)^{2}=I$, which implies for all $n\in\mathbb{N}$ that
$\left(\overrightarrow{\mu}\cdot\overrightarrow{\sigma}\right)^{n}=\begin{cases}I&\text{if $n$ is even}\\ \overrightarrow{\mu}\cdot\overrightarrow{\sigma}&\text{else}\end{cases}.$ (287)
Then we find that
$\exp\!\left(-\beta\overrightarrow{\mu}\cdot\overrightarrow{\sigma}\right)$
$=\sum_{n=0}^{\infty}\frac{\left(-\beta\overrightarrow{\mu}\cdot\overrightarrow{\sigma}\right)^{n}}{n!}$ (288)
$=\sum_{k=0}^{\infty}\frac{\left(-\beta\right)^{2k}\left(\overrightarrow{\mu}\cdot\overrightarrow{\sigma}\right)^{2k}}{\left(2k\right)!}+\sum_{k=0}^{\infty}\frac{\left(-\beta\right)^{2k+1}\left(\overrightarrow{\mu}\cdot\overrightarrow{\sigma}\right)^{2k+1}}{\left(2k+1\right)!}$ (289)
$=\sum_{k=0}^{\infty}\frac{\left(-\beta\right)^{2k}}{\left(2k\right)!}I+\sum_{k=0}^{\infty}\frac{\left(-\beta\right)^{2k+1}}{\left(2k+1\right)!}\overrightarrow{\mu}\cdot\overrightarrow{\sigma}$ (290)
$=\cosh(-\beta)I+\sinh(-\beta)\overrightarrow{\mu}\cdot\overrightarrow{\sigma}$ (291)
$=\cosh(\beta)I-\sinh(\beta)\overrightarrow{\mu}\cdot\overrightarrow{\sigma}.$ (292)
Then
$\operatorname{Tr}\!\left[\exp\!\left(-\beta\overrightarrow{\mu}\cdot\overrightarrow{\sigma}\right)\right]$
$=\operatorname{Tr}[\cosh(\beta)I-\sinh(\beta)\overrightarrow{\mu}\cdot\overrightarrow{\sigma}]$ (293)
$=\cosh(\beta)\operatorname{Tr}[I]-\sinh(\beta)\operatorname{Tr}[\overrightarrow{\mu}\cdot\overrightarrow{\sigma}]$ (294)
$=2\cosh(\beta),$ (295)
which implies that
$\frac{\exp\!\left(-\beta\overrightarrow{\mu}\cdot\overrightarrow{\sigma}\right)}{\operatorname{Tr}\!\left[\exp\!\left(-\beta\overrightarrow{\mu}\cdot\overrightarrow{\sigma}\right)\right]}=\frac{1}{2}\left(I-\tanh(\beta)\overrightarrow{\mu}\cdot\overrightarrow{\sigma}\right)$ (296)
As such, for (283) to be satisfied, the following should hold:
$\frac{1}{2}\left(I+\overrightarrow{r}\cdot\overrightarrow{\sigma}\right)$ $=\frac{1}{2}\left(I-\tanh(\beta)\overrightarrow{\mu}\cdot\overrightarrow{\sigma}\right),$ (297)
which implies that $\overrightarrow{\mu}=-\frac{\overrightarrow{r}}{\left\|\overrightarrow{r}\right\|}$ and $\left\|\overrightarrow{r}\right\|=\tanh(\beta)$, the latter of which is equivalent to $\beta=\operatorname{arctanh}\!\left(\left\|\overrightarrow{r}\right\|\right)$. ∎
Remark 16.
If we would like $\mu$ to be normalized and the temperature to be non-negative, then we could alternatively set $\mu=-\frac{r}{\left\|r\right\|}$ and $\beta=\operatorname{arctanh}\!\left(\left\|r\right\|\right)$, as done in our second proof above, and arrive at the same conclusion. In this way, $\mu$ exclusively represents the direction of the Bloch vector, and the inverse temperature $\beta$ is related only to its norm. This leads to the following corollary, which gives a function that maps $(\mu,\beta)$ to $r$.
Corollary 17.
Let $\mu\in\mathbb{R}^{3}$ be the unit vector (i.e., $\left\|\mu\right\|=1$) of exponential coordinates and $\beta\geq 0$ the inverse temperature such that a qubit state $\rho$ can be written as in (111). Then, by choosing
$r=\tanh(-\beta)\mu,$ (298)
the following equality holds:
$\frac{1}{2}\left(I+r_{x}X+r_{y}Y+r_{z}Z\right)=\\ \frac{1}{Z_{\beta}(\mu)}\exp\!\left(-\beta\left(\mu_{x}X+\mu_{y}Y+\mu_{z}Z\right)\right).$ (299)
Appendix F Proof of Theorem 8
Proof.
We know from (83)–(91) and (96) that the optimal state for (121) has the following form:
$\omega^{*}\equiv\frac{\exp\!\left(-\frac{1}{T}\left(H-T\sum_{i_{1},\ldots,i_{k}}\lambda_{i_{1},\ldots,i_{k}}\overline{\sigma}_{i_{1}}\cdots\overline{\sigma}_{i_{k}}\right)\right)}{\operatorname{Tr}\!\left[\exp\!\left(-\frac{1}{T}\left(H-T\sum_{i_{1},\ldots,i_{k}}\lambda_{i_{1},\ldots,i_{k}}\overline{\sigma}_{i_{1}}\cdots\overline{\sigma}_{i_{k}}\right)\right)\right]},$ (300)
where $\left(\lambda_{i_{1},\ldots,i_{k}}\right)_{i_{1},\ldots,i_{k}\in\left\{0,1,2,3\right\}}$ is a tuple of real coefficients. What remains is to determine the coefficients in $\left(\lambda_{i_{1},\ldots,i_{k}}\right)_{i_{1},\ldots,i_{k}\in\left\{0,1,2,3\right\}}$. To this end, recall that every stabilizer code consists of a Clifford encoding unitary $U$ such that
$U^{{\dagger}}HU$ $=-\sum_{i=1}^{n-k}Z_{i},$ (301)
$U^{{\dagger}}\overline{\sigma}_{i_{1}}\cdots\overline{\sigma}_{i_{k}}U$ $=I^{\otimes n-k}\otimes\sigma_{i_{1}}\otimes\cdots\otimes\sigma_{i_{k}}.$ (302)
Then, after setting
$Z\equiv\operatorname{Tr}\!\left[\exp\!\left(-\frac{1}{T}\left(H-T\sum_{i_{1},\ldots,i_{k}}\lambda_{i_{1},\ldots,i_{k}}\overline{\sigma}_{i_{1}}\cdots\overline{\sigma}_{i_{k}}\right)\right)\right],$ (303)
consider that
$\frac{\exp\!\left(-\frac{1}{T}\left(H-T\sum_{i^{k}}\lambda_{i^{k}}\overline{\sigma}_{i_{1}}\cdots\overline{\sigma}_{i_{k}}\right)\right)}{Z}$
$=\frac{U\exp\!\left(\frac{1}{T}\left(\begin{array}[]{c}\sum_{i=1}^{n-k}Z_{i}+\\ T\sum_{i^{k}}\lambda_{i^{k}}I^{\otimes n-k}\otimes\sigma_{i_{1}}\otimes\cdots\otimes\sigma_{i_{k}}\end{array}\right)\right)U^{{\dagger}}}{Z}$ (306)
$=U\left(\frac{\bigotimes_{i=1}^{n-k}e^{\frac{1}{T}Z}\otimes e^{\sum_{i^{k}}\lambda_{i^{k}}\sigma_{i_{1}}\otimes\cdots\otimes\sigma_{i_{k}}}}{Z}\right)U^{{\dagger}},$ (307)
where we have used the abbreviation $i^{k}\equiv i_{1},\ldots,i_{k}$. Also, we conclude from (306)–(307) that
$Z$ $=\operatorname{Tr}\!\left[\bigotimes_{i=1}^{n-k}e^{\frac{1}{T}Z}\otimes e^{\sum_{i^{k}}\lambda_{i^{k}}\sigma_{i_{1}}\otimes\cdots\otimes\sigma_{i_{k}}}\right]$ (308)
$=\left(\operatorname{Tr}\!\left[e^{\frac{1}{T}Z}\right]\right)^{n-k}\operatorname{Tr}\!\left[e^{\sum_{i^{k}}\lambda_{i^{k}}\sigma_{i_{1}}\otimes\cdots\otimes\sigma_{i_{k}}}\right].$ (309)
Now observe that
$\operatorname{Tr}\!\left[\left(\overline{\sigma}_{i_{1}}\cdots\overline{\sigma}_{i_{k}}\right)\omega^{*}\right]$
$=\operatorname{Tr}\!\left[U^{{\dagger}}\overline{\sigma}_{i_{1}}\cdots\overline{\sigma}_{i_{k}}UU^{{\dagger}}\omega^{*}U\right]$ (310)
$=\operatorname{Tr}\!\left[\left(I^{\otimes n-k}\otimes\sigma_{i_{1}}\otimes\cdots\otimes\sigma_{i_{k}}\right)U^{{\dagger}}\omega^{*}U\right]$ (311)
$=\operatorname{Tr}\!\left[\left(\sigma_{i_{1}}\otimes\cdots\otimes\sigma_{i_{k}}\right)\operatorname{Tr}_{n-k}\!\left[U^{{\dagger}}\omega^{*}U\right]\right].$ (312)
The constraints in (121) and the development in (310)–(312) imply that
$\operatorname{Tr}_{n-k}\!\left[U^{{\dagger}}\omega^{*}U\right]=\rho,$ (313)
where $\rho$ is defined by (118) and (119). However, (307) and (309) imply that
$\operatorname{Tr}_{n-k}\!\left[U^{{\dagger}}\omega^{*}U\right]$
$=\operatorname{Tr}_{n-k}\!\left[\frac{e^{\frac{1}{T}Z_{1}}\otimes\cdots\otimes e^{\frac{1}{T}Z_{n-k}}\otimes e^{\sum_{i^{k}}\lambda_{i^{k}}\sigma_{i_{1}}\otimes\cdots\otimes\sigma_{i_{k}}}}{\left(\operatorname{Tr}[e^{\frac{1}{T}Z}]\right)^{n-k}\operatorname{Tr}\!\left[e^{\sum_{i^{k}}\lambda_{i^{k}}\sigma_{i_{1}}\otimes\cdots\otimes\sigma_{i_{k}}}\right]}\right]$ (314)
$=\frac{e^{\sum_{i_{1},\ldots,i_{k}}\lambda_{i_{1},\ldots,i_{k}}\sigma_{i_{1}}\otimes\cdots\otimes\sigma_{i_{k}}}}{\operatorname{Tr}\!\left[e^{\sum_{i_{1},\ldots,i_{k}}\lambda_{i_{1},\ldots,i_{k}}\sigma_{i_{1}}\otimes\cdots\otimes\sigma_{i_{k}}}\right]}.$ (315)
Given the form in (119) and the equalities in (314)–(315), the only way that the equality in (313) can hold is if $\lambda_{i_{1},\ldots,i_{k}}=\mu_{i_{1},\ldots,i_{k}}$ for all $i_{1},\ldots,i_{k}\in\left\{0,1,2,3\right\}$, thus concluding the proof. ∎
Remark 18.
The development in (306)–(309) indicates that there is a simple method of encoding the state of a stabilizer thermodynamic system. Indeed, the state has the following form:
$U\left(\bigotimes_{i=1}^{n-k}\frac{e^{\frac{1}{T}Z}}{\operatorname{Tr}[e^{\frac{1}{T}Z}]}\otimes\frac{e^{\sum_{i^{k}}\lambda_{i^{k}}\sigma_{i_{1}}\otimes\cdots\otimes\sigma_{i_{k}}}}{\operatorname{Tr}\!\left[e^{\sum_{i^{k}}\lambda_{i^{k}}\sigma_{i_{1}}\otimes\cdots\otimes\sigma_{i_{k}}}\right]}\right)U^{{\dagger}},$ (316)
where we have again used the abbreviation $i^{k}\equiv i_{1},\ldots,i_{k}$. This implies that it can be encoded by preparing the nkn-k ancilla qubits each in temperature-$T$ thermal states of the Hamiltonian Z-Z (which are approximations of the ideal state $|0\rangle\!\langle 0|$) and the $k$ logical qubits in the state
$\frac{e^{\sum_{i^{k}}\lambda_{i^{k}}\sigma_{i_{1}}\otimes\cdots\otimes\sigma_{i_{k}}}}{\operatorname{Tr}\!\left[e^{\sum_{i^{k}}\lambda_{i^{k}}\sigma_{i_{1}}\otimes\cdots\otimes\sigma_{i_{k}}}\right]}.$ (317)
After doing so, one then applies the Clifford encoding unitary $U$ for the stabilizer code.
Appendix G Other simulation results
The technical details of the algorithmic implementations and simulation setup used for the results presented in this appendix are similar to those described in Section VI of the main text. We therefore refer the reader to that section for explanations of implementation-specific aspects of the classical and HQC algorithms.
G.1 One-dimensional quantum Heisenberg model with nearest-neighbor interactions
Figure 8: The figure depicts the logarithm of the error metric in (122) versus the number of iterations, for the first-order classical and HQC algorithms and the second-order classical algorithm for the task of constrained energy minimization for the five-qubit, one-dimensional quantum Heisenberg model with nearest-neighbor interactions and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively. All of the algorithms converge, but the HQC algorithms, shown as the average over five independent runs with shaded regions denoting one standard deviation, require more iterations to converge due to sampling noise inherent in them.
Figure 8 depicts the performance of the first-order classical and HQC algorithms and the second-order classical algorithm for the five-qubit, one-dimensional quantum Heisenberg model with nearest-neighbor interactions and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively.
G.2 Two-dimensional quantum Heisenberg model with nearest-neighbor interactions
Figure 9: The figure depicts the logarithm of the error metric in (122) versus the logarithm of the number of iterations, for the first-order classical and HQC LMPW algorithms for the task of constrained energy minimization for the six-qubit, two-dimensional quantum Heisenberg model with nearest-neighbor interactions and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively. Both algorithms converge, but the HQC algorithm, shown as the average over five independent runs with the shaded region denoting one standard deviation, requires more iterations to converge due to sampling noise inherent in it.
Figure 9 depicts the performance of the first-order classical and HQC algorithms and the second-order classical algorithm for the six-qubit, two-dimensional quantum Heisenberg model with nearest-neighbor interactions and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively. Given the substantial computational cost associated with Hessian estimation in the HQC setting, we did not include the second-order HQC algorithm in this simulation.
G.3 One-to-three-qubit repetition code
The one-to-three qubit repetition code was the earliest discovered quantum error-correcting code [58]. We presented it as a stabilizer thermodynamic system in Example 6.
Figure 10 plots the performance of the LMPW classical first-order algorithm with and without a warm start, showing that warm starting leads to immediate convergence in just one iteration. Figure 11 presents a similar plot for the LMPW HQC first-order algorithm, which likewise benefits from a warm start by converging in a single iteration.
Figure 10: The figure depicts the logarithm of the error metric in (122) versus the number of iterations, for the LMPW classical first-order algorithm for the task of constrained energy minimization for the one-to-three-qubit repetition code. In this case, the plot compares the performance with and without a warm start for the algorithm, as discussed in Section V.1. The warm-started algorithm converges in a single iteration.
Figure 11: The figure depicts the average of the logarithm of the error metric in (122) over five independent runs, plotted against the number of iterations, with shaded regions indicating one standard deviation, for the LMPW HQC first-order algorithm for the task of constrained energy minimization for the one-to-three-qubit repetition code. In this case, the plot compares the performance with and without a warm start for the algorithm, as discussed in Section V.1. The warm-started algorithm converges in a single iteration, and the error bars reflect the standard deviation across the five runs.
G.4 Two-to-four-qubit quantum error-detecting code
The two-to-four qubit quantum error-detecting code encodes two logical qubits into four physical qubits and can detect an arbitrary single-qubit error [61, 62]. Its stabilizer generators are as follows [56, Section 8.1]:
$S_{1}$ $\coloneqq X\otimes X\otimes X\otimes X,$ (318)
$S_{2}$ $\coloneqq Z\otimes Z\otimes Z\otimes Z,$ (319)
and logical operators for it are as follows [56, Eqs. (8.1)–(8.2)]:
$\overline{X}_{1}$ $\coloneqq X\otimes X\otimes I\otimes I,$ (320)
$\overline{Z}_{1}$ $\coloneqq I\otimes Z\otimes I\otimes Z,$ (321)
$\overline{X}_{2}$ $\coloneqq X\otimes I\otimes X\otimes I,$ (322)
$\overline{Z}_{2}$ $\coloneqq I\otimes I\otimes Z\otimes Z.$ (323)
From the above, we deduce that
$\overline{Y}_{1}$ $=i\overline{X}_{1}\overline{Z}_{1}=X\otimes Y\otimes I\otimes Z,$ (324)
$\overline{Y}_{2}$ $=i\overline{X}_{2}\overline{Z}_{2}=X\otimes I\otimes Y\otimes Z,$ (325)
Here, we test the performance of the LMPW algorithms for encoding a Bell state $|\Phi^{+}\rangle\!\langle\Phi^{+}|$, where $|\Phi^{+}\rangle\coloneqq\frac{1}{\sqrt{2}}\left(|00\rangle+|11\rangle\right)$, into the four-qubit error-detecting code. The representation of this state in the Pauli basis is as follows:
$|\Phi^{+}\rangle\!\langle\Phi^{+}|=\frac{1}{4}(I\otimes I+X\otimes X-Y\otimes Y+Z\otimes Z).$ (326)
Thus, we take the non-commuting charges to be all possible logical operators formed from $\overline{X}_{1}$, $\overline{Z}_{1}$, $\overline{X}_{2}$, and $\overline{Z}_{2}$, and the constraints to be as follows:
$\operatorname{Tr}[\overline{X}_{2}\rho]$ $=0,\qquad\operatorname{Tr}[\overline{Y}_{2}\rho]=0,$ (327)
$\operatorname{Tr}[\overline{Z}_{2}\rho]$ $=0,\qquad\operatorname{Tr}[\overline{X}_{1}\rho]=0,$ (328)
$\operatorname{Tr}[\overline{X}_{1}\overline{X}_{2}\rho]$ $=1,\qquad\operatorname{Tr}[\overline{X}_{1}\overline{Y}_{2}\rho]=0,$ (329)
$\operatorname{Tr}[\overline{X}_{1}\overline{Z}_{2}\rho]$ $=0,\qquad\operatorname{Tr}[\overline{Y}_{1}\rho]=0,$ (330)
$\operatorname{Tr}[\overline{Y}_{1}\overline{X}_{2}\rho]$ $=0,\qquad\operatorname{Tr}[\overline{Y}_{1}\overline{Y}_{2}\rho]=-1,$ (331)
$\operatorname{Tr}[\overline{Y}_{1}\overline{Z}_{2}\rho]$ $=0,\qquad\operatorname{Tr}[\overline{Z}_{1}\rho]=0,$ (332)
$\operatorname{Tr}[\overline{Z}_{1}\overline{X}_{2}\rho]$ $=0,\qquad\operatorname{Tr}[\overline{Z}_{1}\overline{Y}_{2}\rho]=0,$ (333)
$\operatorname{Tr}[\overline{Z}_{1}\overline{Z}_{2}\rho]$ $=1.$ (334)
These constraints are chosen in accordance with the representation in (326), in order to ensure that the encoded two-qubit state is indeed a Bell state. Figure 12 plots the performance of the LMPW algorithms for this constrained energy minimization problem.
Figure 12: The figure depicts the logarithm of the error metric in (122) versus the number of iterations, for the first-order classical and HQC algorithms and the second-order classical algorithm for the task of constrained energy minimization for the two-to-four-qubit quantum error-detecting code. All of the algorithms converge, but the HQC algorithms, shown as the average over five independent runs with the shaded region denoting one standard deviation, require more iterations to converge due to sampling noise inherent in them.